### v1

# Structure of the Cayley-Dickson algebras

Abstract
Viewing the Cayley-Dickson process as a graded construction provides a rigorous definition of associativity consisting of three types and the non-associative parts dividing into four types. These simplify the Moufang loop identities and Mal’cev’s identity, which identifies the non-associative Lie algebra structure. This structure is distinct for the first four power-associative algebras and has the same pattern for at least the next three. The structure is identified with 3-cycles and modes that reduce sets of 84 zero divisors to just 7, in most cases, such as for the sedenions, and identifies the subalgebas of the power associative algebras that provide zero divisors thus defining the structure of Cayley-Dickson algebras. The zero divisor and octonion cardinality at all levels is derived. Split algebras are proved to be isomorphic for any power-associative algebra and share the same non-associative structure with different zero divisors. Due to the inclusion of new power-associative algebras into the hypercomplex algebras the terminology ultracomplex numbers is suggested.
Keywords. Octonions, sedenions, Cayley-Dickson algebras, split-sedenions, zero divisors, graded algebras, Moufang loops, Malcev algebra, ultracomplex numbers.
1 Introduction
The Cayley-Dickson construction generates a series of algebras starting with the only four normed division algebras, the reals, complex numbers, quaternions and octonions, denoted ${\mathbb{R}}$, ${\mathbb{C}}$, ${\mathbb{H}}$ and ${\mathbb{O}}$, respectively [1, 2]. After this is an infinite series of power-associative algebras whose properties are analysed herein using a graded notation. These algebras are generally not considered to be graded because pure basis elements have the same properties so permutations are presumed to be isomorphic to each other. Also each level of the construction involves the product of the two Cayley-Dickson algebras at the previous level, which demonstrates that each algebra contains the previous algebra as a subalgebra. This statement and many other properties become obvious once a graded notation is adopted because the Cayley-Dickson process is naturally suited to be graded. The first power-associative algebra is the sedenions, which are not a division algebra due to the existence of zero divisors. The Cayley-Dickson construction includes the sedenions at all higher levels but the associativity properties are different providing structure that is found to be unique for the first four power-associative algebras. Thereafter it is proved that the structure can not become more complicated, just increasing in size so all remaining power-associative algebras have the same structure.
In ungraded form the Cayley-Dickson construction is [3],
$(a,b)(c,d)=(ac+\epsilon d^{*}b,da+bc^{*})\text{,}$
where the conjugate is $(a,b)^{*}=(a^{*},-b)$ and $\epsilon=1$ for the normed composition algebras analysed by Hurwitz [4]. The internal products, $ac$ for instance, are then expanded at the next level and the process repeated generating a polynomial binomial expansion pyramid easily represented by a computer program. This process stops at the reals, ${\mathbb{R}}$. Each new level is generally shown with a new generator operating on the previous level. So quaternions extend $i$ with $j$ providing a new element $k=ij$, which makes the notation ungraded. For octonions, another basis element is introduced, generally $l$, which extends the quaternions in a mixed graded and ungraded notation [4]. A consistent graded notation is obviously needed and the benefits allow the series to be analysed to arbitrary levels, which uncovers different types of associativity that are only exposed in the power-associative algebras.
Deriving an ordering using the graded notation finds that power-associative algebras have three types of associativity, which leads to four types of non-associativity. Octonions only have one symmetric type of non-associativity that makes the three Moufang loop identities identical. Sedenions introduce two more asymmetric non-associative types, which allows for zero divisors and the second power-associative algebra, called the trigintaduonions, includes the final non-associative type which adds more complicated zero divisors. Analysis of the new non-associative types uncovers cyclic forms and modes that categorise zero divisors and applied to the 84 known sedenion zero divisors [5], reduces them to seven distinct triples of basis elements with three cycles and four modes each. The trigintaduonions have 1,260 zero-divisors, which are derived from 147 distinct triples. Zero divisor cardinality is derived as multiples of 84 and copies of ${\mathbb{O}}$ subalgebras are provided. The zero divisors belong to other algebras related to octonions but containing zero divisors hence introducing power associativity into the Cayley-Dickson series of algebras. Analysing these new pseudo-octonion algebras derives the structure of the infinite series. It is found that the pseudo-octonion algebras increase at a fixed rate and the octonions are like remainders. Octonions are the first step in the series existing as a single remainder.
The next section starts by building the Cayley-Dickson construction as two equivalent pyramids using graded and binary notations. Swapping between the two allows the usual composition properties to be proved for all Cayley-Dickson algebras showing that the graded notation is indeed equivalent to the usual doubling of algebras [6]. The structure section derives the four non-associativity types and the eight 3-cycles that present the structure that simplifies the Moufang identities and Malcev algebras. The Malcev algebra defines non-associative Lie algebras and this is used to count subalgebras isomorphic to octonions and some of the power-associative pseudo-algebras. The structure of Cayley-Dickson algebras is derived from the relationship between these subalgebras and octonions. The fourth section defines zero divisors and modes and shows the relationship to the eight 3-cycles. It proceeds to document the seven and 147 prime zero divisors of the first two power-associative algebras and applies the Cycles and Modes theorems to uncover the rest. The final section shows each power-associative algebra has a single split version with the same non-associative structure but different zero divisors. These are provided for the split sedenions.
2 Graded Cayley-Dickson Construction
Definition.
The basis of the Cayley-Dickson algebra at level $n$, called the ungraded dimension, has $2^{n}$ graded elements consisting of grades of size $\binom{n}{k}$ called a $k$-grade. Basis elements are represented as $o_{\alpha}$ where $\alpha$ is a set of digits from ${\mathrm{A}(n)}=\{\emptyset,(1),(12),\ldots,(12\cdots n)\}$, generated by the recurrence relation ${\mathrm{A}(n)}={\mathrm{A}(n-1)}+{\mathrm{A}(n-1)}(n)$ where set multiplication by $(n)$ extends each member of the set by $n$, ${\mathrm{A}(n-1)}(n)=\{(n),(1n),(12n),\ldots,(12\cdots(n-1)n)\}$. This operation is called expansion but multiplying by an existing member will contract some members. Using the notation of Porteous [7] whereby ${\mathbb{N}}_{1}^{n}$ is the set of natural numbers from 1 up to and including $n$, then $\{(12\cdots n)\}(k)=\{(12\cdots(k-1)(k+1)\cdots n)\}$ for $k\in{\mathbb{N}}_{3}^{n-1}$. Note that ${\mathrm{A}(n)}$ includes the empty set, $\emptyset$, or no digit case, to denote the unit graded element, 1, which is the only graded element at level zero and the unit notation is preferred to the use of the empty set, $o=1$. The set brackets are dropped in the graded element, $(\alpha)\rightarrow o_{\alpha}$.
The graded recurrence relation is an example of the binomial theorem, which has the property
$\sum_{k=0}^{n}\binom{n}{k}=2^{n}\text{,}$ (1)
and the basis has a binary expansion described by Pascal’s triangle. This provides a natural ordering of the graded elements shown in the following graded element pyramid to level 3,
$1$
$\{1,o_{1}\}$
$\{1,o_{1},o_{2},o_{12}\}$
$\{1,o_{1},o_{2},o_{12},o_{3},o_{13},o_{23},o_{123}\}$.
Lemma 1.
Any list of distinct graded elements multiplied together via set expansion and contraction generates another distinct graded element if the unit, 1, is excluded.
Proof.
Set expansion and contraction of two distinct members of ${\mathrm{A}(n)}$ is the symmetric difference or the union minus the intersection of the two groups of indices. This is different to the initial two members and this process can continue for any number of distinct graded elements as long as the no digit case is not encountered. Each row of Pascal’s triangle of graded elements contains an even number of each index, excluding level 1, because each index is doubled for the next level of the recurrence relation and there are an even number of existing elements to double with the new index. Hence, excluding the first row, the multiplication of all elements of any row contracts to $\alpha$ being the no digit case. Excluding such cases produces a distinct new graded element. ∎
Definition.
A $k$-grade contains graded elements $o_{\alpha}$ where $\alpha$ has $k$ indices called the length of $\alpha$. The $k=1$ case is called a generator because all graded elements can be generated from these by expansion and contraction. At level $n$ these are $o_{1},o_{2},\cdots,o_{n}$.
Definition.
A pure element excludes the unit basis element, 1. A blade is a pure, non-zero, scaled basis element and a $k$-blade is a member of the $k$-grade. Note that blades may have positive or negative squares and for the remainder this assumed to be ±$\pm 1$, for blades. A blade is positive or negative for scale +1 or -1, respectively.
The Cayley-Dickson construction uses pairs of scalars to represent the graded pyramid.
Definition.
The pyramid element ${a}$, at level $n$, specifies the $n^{\rm th}$ row of a binary pyramid with coefficients $a_{\beta}$, $\beta\in{\mathrm{B}(n)}$, where ${\mathrm{B}(n)}={\mathrm{B}(n-1)}(0)+{\mathrm{B}(n-1)}(1)$ so that ${\mathrm{B}(n)}$ is all binary numbers with $n$ digits from ${\mathbb{N}}_{0}^{1}$ and set multiplication extends the binary numbers in ${\mathrm{B}(n-1)}$. Hence ${\mathrm{B}(n)}$ covers all $2^{n}$ combinations of $n$ digits in ${\mathrm{A}(n)}$ and this is an equivalence relationship to graded elements with coefficients $a_{\beta}=0$ or ±$\pm 1$ because appending a $1$ to $\beta$ at the $n$th position is the same as appending $n$ to the graded recurrence relation. Hence, at level 1, ${a}_{0}=(\pm 1,0)=\pm 1$ and ${a}_{1}=(0,\pm 1)=\pm o_{1}$. To level three this is
${a}$ $=(a_{0},a_{1})$
$=((a_{00},a_{10}),(a_{01},a_{11}))$ (2)
$=(((a_{000},a_{100}),(a_{010},a_{110})),((a_{001},a_{101}),(a_{011},a_{111})))\text{.}$
The Cayley-Dickson process in pyramid form provides the same ordering as the graded elements with the last index corresponding to $o_{n}$, if it is $1$. Some pyramid coefficients at level 2 are $\beta=ij$, $i,j\in{\mathbb{N}}_{0}^{1}$ which corresponds to signed $o_{\beta}=o_{1}\text{,}\>o_{2}\text{,}\>o_{12}$ for $a_{ij}=a_{10}\text{,}\>a_{01}\text{,}\>a_{11}$ in (2), respectively. The four extreme blades at level 3 are
±$\begin{split}\pm o_{1}&=(((0,\pm 1),(0,0)),((0,0),(0,0)))\text{ and}\\ \pm o_{123}&=(((0,0),(0,0)),((0,0),(0,\pm 1)))\text{.}\end{split}$
Definition.
The Cayley-Dickson product at level $i$ for two pure pyramid elements, ${a}_{\alpha}{b}_{\beta}$, with $\alpha,\beta\in{\mathrm{B}(n)}$, with $\alpha_{i}$ being all indices to the left of position $i$ and $\alpha_{{\overline{i}}}$ being all indices to the right of position $i$, is
$\begin{split}(a_{\alpha_{i}{0}\alpha_{\hskip 1.0pt{\overline{i}}}},a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}})(b_{\beta_{i}{0}\beta_{\hskip 1.0pt{\overline{i}}}},b_{\beta_{i}{1}\beta_{\hskip 1.0pt{\overline{i}}}})&=\\ \Big{(}a_{\alpha_{i}{0}\alpha_{\hskip 1.0pt{\overline{i}}}}b_{\beta_{i}{0}\beta_{\hskip 1.0pt{\overline{i}}}}&-\epsilon_{i}b^{*}_{\beta_{i}{1}\beta_{\hskip 1.0pt{\overline{i}}}}a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}},\>b_{\beta_{i}{1}\beta_{\hskip 1.0pt{\overline{i}}}}a_{\alpha_{i}{0}\alpha_{\hskip 1.0pt{\overline{i}}}}+a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}}b^{*}_{\beta_{i}{0}\beta_{\hskip 1.0pt{\overline{i}}}}\Big{)}\text{,}\end{split}$ (3)
where $\epsilon_{i}=\pm 1$ for $i\in{\mathbb{N}}_{1}^{n}$ and the conjugate operator at level $i$ is $a^{*}_{\alpha}0=(a^{*}_{\alpha_{i}{0}\alpha_{\hskip 1.0pt{\overline{i}}}},-a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}})$ and the conjugate operator on the right hand side acts at level $i-1$ and has no effect at level 0. Equation (3) corresponds to a section of the pyramid one pair deep at level $i$. The terms on the right hand side of (3) are called the parallel and cross terms, respectively. The $\alpha_{{\overline{i}}}$ indices are not shown if empty.
The construction uses Pascal’s recurrence relation to create a pyramid, $n$ pairs deep, whereby the first division repeats the structure on of the left half on the right but multiplied by $o_{n}$. For $n=1$, so that $\alpha,\beta\in{\mathrm{B}(0)}$, the no index case, then
$(a_{0},a_{1})(b_{0},b_{1})=(a_{0}b_{0}-\epsilon_{1}b_{1}a_{1},b_{1}a_{0}+a_{1}b_{0})$ (4)
since $a_{0},a_{1},b_{0},b_{1}\in{\mathbb{R}}$ and the conjugate has no effect at level zero.
For $n\geq 2$, showing just the first pair in each half, with $\alpha\in{\mathrm{B}(n)}$ and $\beta=00\cdots 0$ of length $n-2$, then
${a}_{\alpha}=\Big{(}\overbrace{\underbrace{\big{(}(\cdots(}_{n-2}(a_{\beta 00},a_{\beta 10}),\ldots)\big{)}}^{n-1\rm\;pairs},\overbrace{\underbrace{\big{(}(\cdots(}_{n-2}(a_{\beta 01},a_{\beta 11}),\ldots)\big{)}}^{n-1\rm\;pairs}\Big{)}\text{.}$ (5)
For $n=2$, again with $\alpha,\beta\in{\mathrm{B}(1)}$ of lengths 1,
$\begin{split}&(a_{0},a_{1})(b_{0},b_{1})\\ &=\big{(}(a_{00},a_{10}),(a_{01},a_{11}))((b_{00},b_{10}),(b_{01},b_{11})\big{)}\\ &=\big{(}(a_{00},a_{10})(b_{00},b_{10})-\epsilon_{2}(b_{01},b_{11})^{*}(a_{01},a_{11}),\\ &\enspace\enspace\enspace\enspace(b_{01},b_{11})(a_{00},a_{10})+(a_{01},a_{11})(b_{00},b_{10})^{*}\big{)}\\ &=\big{(}(a_{00}b_{00}-\epsilon_{1}b_{10}a_{10},b_{10}a_{00}+a_{10}b_{10})-\epsilon_{2}(b_{01}a_{01}+\epsilon_{1}a_{11}b_{11},b_{01}a_{11}-b_{11}a_{01}),\\ &\enspace\enspace\enspace\enspace\big{(}b_{01}a_{00}-\epsilon_{1}a_{10}b_{11},a_{10}b_{01}+b_{11}a_{00})+(a_{01}b_{00}+\epsilon_{1}b_{10}a_{11},-b_{10}a_{01}+a_{11}b_{00})\big{)}\\ &=\big{(}(a_{00}b_{00}-\epsilon_{1}b_{10}a_{10}-\epsilon_{2}b_{01}a_{01}-\epsilon_{1}\epsilon_{2}a_{11}b_{11},b_{10}a_{00}+a_{10}b_{00}-\epsilon_{2}b_{01}a_{11}+\epsilon_{2}b_{11}a_{01}),\\ &\enspace\enspace\enspace\enspace\big{(}b_{01}a_{00}-\epsilon_{1}a_{10}b_{11}+a_{01}b_{00}+\epsilon_{1}b_{10}a_{11},a_{10}b_{01}+b_{11}a_{00}-b_{10}a_{01}+a_{11}b_{00})\big{)}\text{.}\end{split}$ (6)
Lemma 2.
For blade $o_{i_{1}i_{2}\ldots i_{k}}$ then $o_{i_{1}i_{2}\ldots i_{k}}^{2}=-\epsilon_{i_{1}}\epsilon_{i_{2}}\ldots\epsilon_{k}$, for $i_{1}i_{2}\ldots i_{k},k\in{\mathbb{N}}_{1}^{n}$.
Proof.
From (6) or (3) with $\alpha=\beta$ and $\beta\in{\mathrm{B}(1)}$ then only parallel terms exist at level two,
$\begin{split}\sum_{\beta}(a_{\beta 0},a_{\beta 1})(b_{\beta 0},b_{\beta 1})&=((a_{00}b_{00}-\epsilon_{1}b_{10}a_{10}-\epsilon_{2}b_{01}a_{01}-\epsilon_{1}\epsilon_{2}a_{11}b_{11},0),\\[-12.91663pt] &\enspace\enspace\enspace\enspace(b_{01}a_{00}-\epsilon_{1}a_{10}b_{11}+a_{01}b_{00}+\epsilon_{1}b_{10}a_{11},0))\text{.}\end{split}$
Hence, by symmetry and embedding in (5), the following apply to levels of the pyramid after one,
$\begin{split}o_{1}o_{1}&=((-\epsilon_{1}b_{10}a_{10},0),(0,0))=-\epsilon_{1}\text{,}\\ o_{2}o_{2}&=((-\epsilon_{2}b_{01}a_{01},0),(0,0))=-\epsilon_{2}\text{ and}\\ o_{12}o_{12}&=((-\epsilon_{1}\epsilon_{2}a_{11}b_{11},0),(0,0))=-\epsilon_{1}\epsilon_{2}\text{.}\end{split}$
This is generalised to a $k$-blade, $a_{\alpha}$, $\alpha\in{\mathrm{B}(n)}$, with $\alpha$ containing ones at positions $i_{1},i_{2},\ldots,i_{k}\in{\mathbb{N}}_{1}^{n}$, $k\in{\mathbb{N}}_{1}^{n}$, and zeros elsewhere. From (3), if $i_{k}\neq n$ then the zeros at the end of $\alpha$ will act as $a_{\alpha}^{2}=a_{\alpha}b_{\alpha}$ until the rightmost index, $i_{k}=i$, is reached. Then
$a_{\alpha}b_{\alpha}=a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}}b_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}}=(-\epsilon_{i}b^{*}_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}}a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}},0)$
where
[…]
The first case is positive because the original conjugation negated the $o_{i}$ element and a new conjugate for the $i-1$ level was introduced whereas for $b^{*}_{\alpha_{i-1}{0}\alpha_{\hskip 1.0pt{\overline{i-1}}}}$ the conjugation is passed to the next level without effect. Continuing to the following levels has no effect on the original sign and together will introduce $\epsilon_{j}$ from $a^{2}_{{\alpha_{j}{1}\alpha_{\hskip 1.0pt{\overline{j}}}}}\forall j\in i_{1},i_{2},\ldots,i_{k-1}$. This means $o_{i_{1}i_{2}\ldots i_{k}}^{2}=-\epsilon_{i_{1}}\epsilon_{i_{2}}\cdots\epsilon_{i_{k}}$ which is assumed true and by induction $a^{2}_{\alpha 1}$ is considered for $\alpha$ having ones in positions $i_{1}i_{2}\ldots i_{k}$ and zeros elsewhere. Then
$a_{\alpha 1}b_{\alpha 1}=(-\epsilon_{n+1}b^{*}_{\alpha 1}a_{\alpha 1},0)\text{,}$
hence $o_{i_{1}i_{2}\ldots i_{k}(n+1)}^{2}=-\epsilon_{i_{1}}\epsilon_{i_{2}}\ldots\epsilon_{i_{k}}\epsilon_{n+1}$. ∎
Lemma 3.
For $\delta\in{\mathrm{A}(n-1)}$ not all zeros then $o_{\delta}o_{n}=o_{\delta n}$ and $o_{n}o_{\delta}=-o_{\delta n}$.
Proof.
At level $n=1$, $1o_{1}=(1,0)(0,1)=(0,b_{1}a_{0})=o_{1}$. At $n=2$,
$o_{1}o_{2}=((0,1),(0,0))((0,0),(1,0))=((0,0),(0,a_{10}b_{01}))=o_{12}\text{.}$
For the reverse product, at level $n=1$, $o_{1}1=(0,1)(1,0)=(0,b_{1}a_{0})=o_{1}$. At $n=2$,
$o_{2}o_{1}=((0,0),(1,0))((0,1),(0,0))=((0,0),(0,a_{01}b^{*}_{10}))=-o_{12}\text{.}$
The symmetry of the $o_{n}$ pyramid used here for any $o_{\delta}\equiv a_{\alpha 0}$ is that the first term on the right hand side of the pyramid is a cross term and all the rest are parallel terms,
$a_{\alpha 0}b_{\beta 1}=(0,a_{\alpha 0}b_{\beta 1})=\begin{cases}\big{(}(0,0),(0,a_{\alpha_{n-1}10}b^{*}_{\beta_{n-1}01})\big{)}&\text{if $o_{\delta}$ contains $o_{n-1}$ and}\\ \big{(}(0,0),(a_{\alpha_{n-1}00}b_{\beta_{n-1}01},0)\big{)}&\text{otherwise.}\end{cases}$
Since conjugation has no effect on $b_{\beta_{n-1}0}$ then extrapolating to level 0 gives a positive sign, $o_{\delta}o_{n}=o_{\delta n}$. The second case, $o_{n}o_{\delta}$, has the same symmetry, with the conjugate having no effect other than the minus sign introduced at level n,
$a_{\beta 1}b_{\alpha 0}=(0,a_{\beta 1}b^{*}_{\alpha 0})=\begin{cases}\big{(}(0,0),(0,-b_{\alpha_{n-1}10}a^{*}_{\beta_{n-1}01})\big{)}&\text{if $o_{\delta}$ contains $o_{n-1}$ and}\\ \big{(}(0,0),(a_{\beta_{n-1}01}b_{\alpha_{n-1}00},0)\big{)}&\text{otherwise}\end{cases}$
hence $o_{n}o_{\delta}=-o_{\delta n}$. ∎
Theorem 1 (Cayley-Dickson algebras).
The Cayley-Dickson construction to any level, $n$, generates an algebra with pure elements, $x$, $y$ satisfying $xy=-yx$ if $x\neq y$ and, if $\epsilon_{i}=1$, $\forall i\in{\mathbb{N}}_{1}^{n}$ and $x=y$, then $xy=-1$. The unit, 1, has no effect under multiplication.
Proof.
From Lemma 1, product expansion and contraction of blades shows that the generators of the algebra under multiplication are an injection and repeated multiplication becomes a surjection thus generating the scaled basis of the algebra. Lemma 2 shows that for $\epsilon_{i}=1$, $\forall i\in{\mathbb{N}}_{1}^{n}$ then $o_{\alpha}^{2}=-1$, $\forall\alpha\in{\mathrm{A}(n)}$. For the unit, $1=b_{\beta}$ where $\beta\in{\mathrm{B}(n)}$ is all zeros, the product from (3) at level $i>1$ is seen to be an identity for itself and pure elements,
$(a_{\alpha_{i}{0}\alpha_{\hskip 1.0pt{\overline{i}}}},a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}})(b_{\beta},0)$ $=\left(a_{\alpha_{i}{0}\alpha_{\hskip 1.0pt{\overline{i}}}}b_{\beta},a_{\alpha_{i}{1}\alpha_{\hskip 1.0pt{\overline{i}}}}b^{*}_{\beta}\right)$
$=\big{(}(a_{\alpha_{i-1}{0}\alpha_{\hskip 1.0pt{\overline{i-1}}}},0)(b_{\beta},0),(a_{\alpha_{i-1}{1}\alpha_{\hskip 1.0pt{\overline{i-1}}}},0)(b_{\beta},0)\big{)}\text{.}$
The expansion terminates at level 0 with $a_{\alpha}b_{\beta}=a_{\alpha}$ and is symmetric so $a_{\beta}b_{\alpha}=b_{\alpha}$ also.
Anticommutation is proved by induction starting with Lemma 3 for $o_{1}o_{2}=-o_{2}o_{1}$. This is the $n=2$ level but is provided as the basis of induction in the proof of the lemma. By the symmetry of the construction, this result applies at all levels $n>2$ for these pure elements. Assuming the $n-1$ product is true, $a_{\alpha 0}b_{\beta 0}=-b_{\beta 0}a_{\alpha 0}$, for $\alpha,\beta\in{\mathrm{B}(n-1)}$, not all zeros, and $\alpha\neq\beta$, then multiplying by $o_{n}$ produces a pyramid whereby the left-hand half does not contain $o_{n}$ and the right-hand half reflects this structure but includes $o_{n}$ for all elements. Proving $xy=-yx$ for the $o_{n}$ case is divided into three equations which need to be proved; $o_{n}a_{\alpha 0}=-a_{\alpha 0}o_{n}$, $a_{\alpha 1}b_{\beta 0}=-b_{\beta 0}a_{\alpha 1}$ and $a_{\alpha 1}b_{\beta 1}=-b_{\beta 1}a_{\alpha 1}$, for $\alpha,\beta$ not all zeros. Lemma 3 has already shown the first case and found that $a^{*}_{\alpha}=-a_{\alpha}$ since $\alpha$ is not all zeros. Hence, the second case with $\beta$ not all zeros, is
$\begin{split}{a}_{\alpha 1}{b}_{\beta 0}&=(0,a_{\alpha 1})(b_{\beta 0},0)=(0,a_{\alpha 1}b_{\beta 0}^{*})\text{ and}\\ {b}_{\beta 0}{a}_{\alpha 1}&=(b_{\beta 0},0)(0,a_{\alpha 1})=(0,a_{\alpha 1}b_{\beta 0})=-{a}_{\alpha 1}b_{\beta 0}\text{.}\end{split}$
Finally, induction and symmetry of the pyramid is used on $a_{\alpha 1}b_{\beta 1}+b_{\beta 1}a_{\alpha 1}=0$,
$\begin{split}{a}_{\alpha 1}{b}_{\beta 1}&=(0,a_{\alpha 1})(0,b_{\beta 1})=(-b_{\beta 1}^{*}a_{\alpha 1},0)\text{ and}\\ {b}_{\beta 1}{a}_{\alpha 1}&=(0,b_{\beta 1})(0,a_{\alpha 1})=(-a_{\alpha 1}^{*}b_{\beta 1},0)\text{.}\end{split}$ (7)
These anticommutator terms occur in the same position on the right-hand half as the terms without $o_{n}$ on the left-hand half of the pyramid. After the first pyramid multiplication they have identical paths through the each side of the pyramid apart from the minus signs and the conjugate. The conjugate descends the path until the first 1 in $\alpha$ from the right and changes the sign and disappears for the rest of the path, $a_{\alpha 1}^{*}=-a_{\alpha 1}$ and similarly, $a_{\beta 1}^{*}=-a_{\beta 1}$. Hence the minus signs are cancelled in (7) and so using symmetry and arguing by induction with $a_{\alpha 0}b_{\beta 0}-b_{\beta 0}a_{\alpha 0}=0$ true at level 2 and assumed true at level $n-1$, then
$-b_{\beta 1}^{*}a_{\alpha 1}-a_{\alpha 1}^{*}b_{\beta 1}=a_{\alpha 0}b_{\beta 0}+b_{\beta 0}a_{\alpha 0}=0\text{.}$
Corollary.
The first four levels of blades with $\epsilon_{i}=1\;\forall i\in{\mathbb{N}}_{1}^{4}$ are isomorphic to complex numbers, ${\mathbb{C}}$, quaternions, ${\mathbb{H}}$, octonions, ${\mathbb{O}}$, and sedenions, ${\mathbb{S}}$, respectively.
Proof.
Theorem 1 showed that blades generate Cayley-Dickson algebras under the product rules (3). It is obvious that the first generator, $o_{1}$, at level 1 is isomorphic to the imaginary of complex numbers. At level two, the usual representation of quaternions also have the cyclic product rules, $ij=k$, $jk=i$ and $ki=j$. Starting by calculating the cyclic rules in full
$\begin{split}o_{1}o_{2}&=((0,1),(0,0))((0,0),(1,0))=((0,0),(0,1)(1,0))\\ &=((0,0),(0,1^{*}))=o_{12}\text{,}\\ o_{2}o_{12}&=((0,0),(1,0))((0,0),(0,1))=(-(0,1)^{*}(1,0),(0,0))\\ &=((0,1),(0,0))=o_{1}\ \text{ and}\\ o_{12}o_{1}&=((0,0),(0,1))((0,1),(0,0))=((0,0),(0,1)(0,1)^{*})\\ &=((0,0),-(-1,0))=o_{2}\text{.}\end{split}$ (8)
Thus level 2 is isomorphic to ${\mathbb{H}}$. At level 3 there are seven blades, $\{o_{1},o_{2},o_{12},o_{3},o_{23},o_{13},o_{123}\}$, which matches the number of octonion basis elements and it is obvious that the quaternions are embedded in the first three pure terms, which is a feature of ${\mathbb{O}}$. This suffices to provide the isomorphism. At level 4 there are 15 blades matching the sedenions and the octonions are embedded in the first seven blades. This is a requirement of sedenions and will suffice to prove the isomorphism, given Theorem 1. ∎
The alternative ring properties of Cayley-Dickson algebras that are essential to simplify non-associative expressions are now proved [8].
Lemma 4.
For $x$ and $y$ distinct blades with $\epsilon_{i}=1$, $i\in{\mathbb{N}}_{1}^{n}$, then $yxx=-y$ and $x(xy)=-y$.
Proof.
The later statement is true if the former is true, $x(xy)=-(xy)x=yxx$, if $y\neq x$ and obviously $x(xx)=-x$ otherwise. The proof of the former proceeds via induction. It works for ${\mathbb{H}}$ which needs to be checked for the six combinations of two from three elements using the cyclic product rules (8),
$o_{1}o_{2}o_{2}$ $=o_{12}o_{2}=-o_{2}o_{12}=-o_{1}\text{,}$
$o_{1}o_{12}o_{12}$ $=-o_{2}o_{12}=o_{12}o_{2}=-o_{1}\text{,}$
$o_{2}o_{1}o_{1}$ $=-o_{1}o_{2}o_{1}=-o_{12}o_{1}=-o_{2}\text{,}$
$o_{2}o_{12}o_{12}$ $=-o_{12}o_{2}o_{12}=o_{1}o_{12}=-o_{2}\text{,}$
$o_{12}o_{1}o_{1}$ $=-o_{1}o_{12}o_{1}=o_{2}o_{1}=-o_{12}\text{ and}$
$o_{12}o_{2}o_{2}$ $=-o_{1}o_{2}=-o_{12}\text{.}$
Assuming ${a}_{\alpha 0}{b}_{\beta 0}{b}_{\beta 0}={b}_{\beta 0}({b}_{\beta 0}{a}_{\alpha 0})=-{a}_{\alpha 0}$, $\alpha,\beta\in{\mathrm{B}(n-1)}$ not all zeros, then by induction only the following conditions need to be proved, which for the second two equations again use the symmetry argument that after the first step in the pyramid expansion the left-right symmetry means the same path is followed on both sides,
${a}_{\alpha 1}{b}_{\beta 0}{b}_{\beta 0}$ $=(0,a_{\alpha 1})(b_{\beta 0},0)(b_{\beta 0},0)$
$=(0,a_{\alpha 1}b_{\beta 0}^{*})(b_{\beta 0},0)$
$=(0,a_{\alpha 1}b_{\beta 0}^{*}b_{\beta 0})$
$=-{a}_{\alpha 1}\text{,}$
${a}_{\alpha 0}{b}_{\beta 1}{b}_{\beta 1}$ $=(a_{\alpha 0},0)(0,b_{\beta 1})(0,b_{\beta 1})$
$=(0,b_{\beta 1}a_{\alpha 0})(0,b_{\beta 1})$
$=(-b_{\beta 1}^{*}(b_{\beta 1}a_{\alpha 0}),0)$
$=(b_{\beta 0}(b_{\beta 0}a_{\alpha 0}),0)$
$=-{a}_{\alpha 0}\quad\text{ and }$
${a}_{\alpha 1}{b}_{\beta 1}{b}_{\beta 1}$ $=(0,a_{\alpha 1})(0,b_{\beta 1})(0,b_{\beta 1})$
$=(-b_{\beta 1}^{*}a_{\alpha 1},0)(0,b_{\beta 1})$
$=(0,-b_{\beta 1}(b_{\beta 1}^{*}a_{\alpha 1}),0)$
$=(0,b_{\beta 0}(b_{\beta 0}a_{\alpha 1}),0)$
$=(0,-a_{\alpha 1})$
$=-{a}_{\alpha 1}\text{.}$
This leaves $xo_{n}o_{n}$ where $\beta$ is all zeros in the last two equations. In that case the conjugate has no effect after the first step and $-b_{\beta 1}^{*}(b_{\beta 1}x)=-b_{\beta 1}(b_{\beta 1}^{*}x)=-x$. Hence ${a}_{\alpha}{b}_{\beta}{b}_{\beta}=-{a}_{\alpha}$ for all $\alpha,\beta\in{\mathrm{B}(n)}$. Note that while $xyzz=-xy$ it is not true that $x(xyz)=-yz$ for a distinct blade $z$. For example, $o_{1}(o_{1}o_{2}o_{3})=o_{23}$. ∎
Definition.
The associator is defined, using left-expansion, as
$[a,b,c]=(ab)c-a(bc)=abc-a(bc)\text{,}$
which is zero for ${\mathbb{H}}$ because quaternions are associative [3] and have one ordered pure triple. This is not the case for octonions, ${\mathbb{O}}$, which have $\binom{7}{3}=35$ ordered pure triples. Of these 28 have a non-zero associator and seven are associative. The latter triples are
$\begin{split}&(o_{1},o_{2},o_{12}),(o_{2},o_{3},o_{23}),(o_{1},o_{3},o_{13}),(o_{1},o_{23},o_{123}),\\ &(o_{2},o_{13},o_{123}),(o_{12},o_{3},o_{123}),(o_{12},o_{13},o_{23})\text{.}\end{split}$ (9)
All of the triples in (LABEL:eqn:qa7) are quaternion-like, either isomorphic or anti-isomorphic to quaternions, with cyclic product rules and the first one is the quaternion subset from level 2. The next two have the same structure and would be called quaternions if the basis where reordered. The last triple provides rotations between the three generators, which is a feature of quaternions if the generators represent three dimensional vector space, ${\mathbb{R}}^{3}$. The three remaining triples can not be interpreted in this way but the graded product of the three blades in all seven triples is seen to be a scalar, either ±$\pm 1$. The remaining triples will be shown to be non-associative. For example, Lemma 2 has $o_{1}o_{2}o_{3}=o_{123}$ whereas expanding the pyramid for $o_{1}(o_{2}o_{3})$ gives the term $-a_{100}b_{011}$ in the $o_{123}$ position so that $o_{1}(o_{2}o_{3})=-o_{123}$, hence the associator is non-zero and this triple is non-associative, $[o_{1},o_{2},o_{3}]=2o_{123}$.
The names quaternion and octonion seem to be derived from book binding techniques involving a quire of four or eight folios of paper, respectively. Traditionally, an ancient book or codex would use separate sheets or bifolio with a single fold stacked and bound to the binding or stub. This was superseded with folding a single large sheet several times to make the quire. The type of fold starts with a bifolio, which is double the number of folios. Two bifolia have quarto fold type, four has Octavo and eight has extodecimo type, which covers more than eight folios [9]. Accordingly, the prefix ultra is introduced herein to designate all power-associative Cayley-Dickson algebras. Then, the complex numbers are followed by the hypercomplex numbers, ${\mathbb{O}}$ and ${\mathbb{H}}$, and the following power-associative algebras would be known as the ultracomplex numbers.
Definition.
The Cayley-Dickson series for power-associative algebras after level $3$ with $\epsilon_{i}=1$ for $i\in{\mathbb{N}}_{1}^{n}$ is designated ${\mathbb{U}_{m}}$, $m=n-3$, which refers to the ultronion algebras. Thus the normed algebras, ${\mathbb{R}}$, ${\mathbb{C}}$, ${\mathbb{H}}$ and ${\mathbb{O}}$, are not included in the ${\mathbb{U}_{m}}$ series and the series starts with the sedenions, ${\mathbb{S}}\equiv{\mathbb{U}_{1}}$. The split-algebras are denoted ${\rm\widetilde{\mathbb{U}}_{m}}$ and later it is shown they are isomorphic to each other, for any dimension, for any positive number of $\epsilon_{i}=-1$, $i\in{\mathbb{N}}_{1}^{n}$. The proof of uniqueness of ${\rm\widetilde{\mathbb{U}}_{n}}$ also applies to split versions of ${\mathbb{C}}$ (Lorentz numbers), ${\mathbb{H}}$ and ${\mathbb{O}}$.
3 Structure Theorems
To concentrate on non-associative triple products the term triad is used to provide distinct triples that avoid scalars and repeated elements with trivial associativity.
Definition.
There are $N^{3}$ scaled triple products $bcd$ and $\binom{N}{3}$ naturally ordered triples with blades $b<c<d$ where $N=2^{n}-1$ and $n$ is the number of generators. The ordered triples are the same as removing all pairs and triples of the same element from the $N^{3}$ set and dividing by the number of permutations, $6$. A triad is defined as ordered triples containing distinct blades whereby $b<c<d$ so that $bcd$ is non-zero. Hence basis element $a=bcd$ can be defined for triads and $a$ is distinct from $b,c,d$, which follows from Lemma 1. Of course, there are 5 unordered triples for each triad.
Theorem 2 (Associativity theorem).
The $24$ permutations of triad products generate three types of associativity, called Types 1, 2 and 3. These are shown in Table 1 where equivalence of associators means both are zero or both are non-zero, noting that only 12 permutations for ordered $b,c,d$ and $a=bcd$ are required because $[a,b,c]=-[c,b,a]$.
Table 1: Triad Associativity Types
Associativity Equivalence Associativity
$[b,a,c]\approx[b,d,c]\approx[a,b,d]\approx[a,c,d]$ Type 1
$[a,b,c]\approx[b,c,d]\approx[b,a,d]\approx[a,d,c]$ Type 2
$[a,c,b]\approx[c,b,d]\approx[a,d,b]\approx[c,a,d]$ Type 3
Proof.
Excluding complete associativity for triads where all associativity types are zero means that $a(bc)=-(bc)a=cba$ if $a$ is not a scaling of $bc$ and similarly for $abc=(ab)c=c(ba)$. Hence $[a,b,c]=abc-a(bc)=c(ba)-(cb)a=-[c,b,a]$. This is also true for associative triple products, where $a$ is a scalar, because both sides are zero. Now, if $b,c,d$ are not associative, the relation $a=bcd$ means $ad^{-1}=bc$ or $d=-a(bc)^{-1}=-bca/b^{2}/c^{2}$, noting that $(bc)^{2}=-b^{2}c^{2}$. First, Type 2 and Type 3 associativity are analysed.
Assuming Type 2 associativity $[a,b,c]=abc-a(bc)=0$ or $[c,b,a]=cba-c(ba)=0$, then substituting $d$ into the other three associators in the second row in Table 1, scaled by $(bc)^{2}$, gives
${[}b,c,d](bc)^{2}$ $=bc(bca)-b(c(bca))=-(bca)(bc)-(bca)cb=-(bc)a(bc)+c(ba)cb$
$=a(bc)(bc)-baccb=a(bc)^{2}+abbc^{2}=-ab^{2}c^{2}+ab^{2}c^{2}=0\text{,}$
${[}b,a,d](bc)^{2}$ $=ba(bca)-b(a(bca))=-ba(c(ba))+b(bc)a^{2}=c(ba)(ba)+cbba^{2}=0\text{\; and}$
$[a,d,c](bc)^{2}$ $=a(bca)c-a(bcac)=-(bca)ac+a(c(ba)c)=-bcca^{2}+(bacc)a=0\text{.}$
Assuming Type 3 associativity $[a,c,b]=acb-a(cb)=0$ or $[b,c,a]=bca-b(ca)=0$, then substituting $d$ into the other three associators in the third row of Table 1, scaled by $(bc)^{2}$, gives
$[c,b,d](bc)^{2}$ $=cb(bca)-c(b(bca))=-b(ca)(cb)-bcabc=cab(cb)-b(ca)bc$
$=-a(cb)(cb)+(ca)bbc=a-cac=0$
${[}a,d,b](bc)^{2}$ $=a(bca)b-a(bcab)=-bcaab+bcaba=bcb+b(ca)ba=0\text{\; and}$
${[}c,a,d](bc)^{2}$ $=ca(bca)-c(a(bca))=-bca(ca)-bcaac=-b(ca)(ca)+bcc=0\text{.}$
Assuming Type 1 associativity $[b,a,c]=bac-b(ac)=0$ or $[c,a,b]=cab-c(ab)=0$, then substituting $d$ into the other three associators in the first row of Table 1, scaled by $(bc)^{2}$, gives
${[}b,d,c](bc)^{2}$ $=b(bca)c-b(bcac)=-bcabc+bcacb\text{,}$
$[a,b,d](bc)^{2}$ $=ab(bca)-a(b(bca))=(bca)(ba)-bcaba=-cba(ba)+cbaba\text{ and}$
${[}a,c,d](bc)^{2}$ $=ac(bca)-a(c(bca))=-bca(ac)-bcaca\text{.}$
If Type 2 associativity is also assumed then $cba=c(ba)$ and, also again using $[c,a,b]=0$, the three equations become
$-bcabc+bcacb$ $=c(ba)bc-c(ba)cb=-b(ac)bc+baccb=acbbc+abb=0\text{,}$
$-cba(ba)+cbaba$ $=c(ba)(ba)-c(ba)ba=-c+bac(ab)=-c+b(ac)ba=-c+cabba=0\text{ and}$
$-bca(ac)-bcaca$ $=c(ba)(ac)+c(ba)ca=-bac(ac)-bacca=-b(ac)(ac)-b=0\text{.}$
If Type 3 associativity is also assumed with Type 1 then $bca=b(ca)$ or $cba=-b(ca)$ and, also again using $[c,a,b]=0$, the three equations become
$-bcabc+bcacb$ $=-b(ca)bc+b(ca)cb=-acbbc-cabcb=-a-c(ab)cb=-a+abccb=0\text{,}$
$-cba(ba)+cbaba$ $=b(ca)(ba)-b(ca)ba=-cab(ba)+cabba=c(ab)(ab)+c=0\text{ and}$
$-bca(ac)-bcaca$ $=b(ca)(ca)-b(ca)ca=-b+cabca=-b+c(ab)ca=-b+bacca=0\text{.}$
Hence Types 2 and 3 associativity are non-associatively equivalent in there own right but Type 1 needs either Type 2 or Type 3 for equivalence. ∎
Definition.
The triple associator, $T(b,c,d)$, is introduced to cover associativity Types 1, 2 and 3 to allow for an unambiguous definition of associativity,
$T(b,c,d)=[b,d,c]-[d,c,b]+[c,b,d]\text{.}$
Definition.
A triad, $b,c,d$, is associative if $T(b,c,d)=0$ and non-associative if $T(b,c,d)\neq 0$. This may seem like over-kill but will soon be seen to be a necessary condition.
Artin’s theorem [3] is stated only for associative subalgebras of octonions but can be generalised and then extended to find all subalgebras of any Cayley-Dickson algebra.
Theorem 3 (Generalised Artin theorem).
Any two distinct blades of ${\mathbb{H}}$, ${\mathbb{O}}$ or ${\mathbb{U}_{m}}$, $m>0$, generate an algebra isomorphic to quaternions which is thus associative.
Proof.
From Lemma 1, multiplying any two distinct blades generates a third distinct blade. Since the generated blade is the symmetric difference then multiplying this third element by one of the first two will recover the other of the first two. All such products are closed under multiplication and forms a Cayley-Dickson algebra, by Theorem 1, because distinct blades anti-commute and have negative square in this case. For any two blades, $b$ and $c$ then $bc=d$, $cd=c(bc)=-bcc=b$ and $db=-cbb=c$, by Lemma 4. These are the cyclic, quaternion-like product rules and thus the triad $b$, $c$ and $bc$ is isomorphic to ${\mathbb{H}}$. The triple associator shows associativity for the triad,
$\begin{split}T(b,c,bc)&=b(bc)c-b(bcc)-bc(cb)+bc(cb)+cb(bc)-c(b(bc))\\ &=-bcbc+bb-bc(bc)+c(bcb)\\ &=+1-1+1-1\\ &=0\text{.}\\[-17.22217pt] \end{split}$
Theorem 4 (Ultronion generation).
Any ultacomplex algebra, ${\mathbb{U}_{m}}$, $m>0$, can be generated by $n=m+3$ distinct blades, $o_{\alpha_{k}},\forall\>k\in{\mathbb{N}}_{1}^{n}$, with indices $\alpha_{k}$ that include all $n$ indices an odd number of times over all $k$.
Proof.
Triad $b,c,d$ (ie three distinct blades), with $a=bcd$ non-scalar, is non-associative,
$\begin{split}T(b,c,d)&=bdc-b(dc)-dcb+d(cb)+cbd-c(bd)\\ &=bdc+dcb-dcb+bcd-bcd+bdc\\ &=2bdc\\ &=\pm 2a\text{.}\end{split}$ (10)
Combinations of $b,c,d$ give seven distinct blades by Lemma 1, $bc,bd,cd,bcd$, which are isomorphic to octonions because they satisfy the rules of ${\mathbb{U}_{n}}$ and these are the same as ${\mathbb{O}}$ by Theorem 1. Any $n$ such terms similarly generate $12^{n}-$ 1 distinct terms if, by Lemma 1, the symmetric difference of their grades do not become the empty set, which is guaranteed by the condition that each generator is represented an odd number of times in the indices. Hence they are isomorphic to ${\mathbb{U}_{n-3}}$, by Theorem 1. ∎
The generation theorem disguises the complexity of the algebras since it avoids generating algebras from arbitrary triads at higher levels. This will take some effort and starts by analysing the asymmetry of non-associativity.
Theorem 5 (Non-associativity theorem).
If a triad is not associative then either all associativity types are non-zero or only one is non-zero, the other two being zero.
Proof.
From Table 1, for triad $b,c,d$, Type 1 associativity has $bdc=b(dc)$, Type 2 associativity has $bcd=b(cd)$ and Type 3 associativity has $cbd=c(bd)$. Starting with Type 1 and substituting Type 3 then Type 2 associativity, is
$bdc=-c(db)=-cbd=bcd=-b(dc)\text{,}$
which contradicts Type 1 associativity if Types 2 and 3 are assumed. Similarly, substituting the reverse of Type 3 then Type 1 into Type 2,
$bcd=-d(bc)=-dbc=bdc=b(dc)=-b(cd)\text{,}$
which contradicts Type 2 associativity. Finally substituting the reverse of both Type 2 and Type 1 into Type 3,
$cbd=-d(cb)=-dcb=cdb=c(db)=-c(bd)\text{,}$
which contradicts Type 3 associativity. This proves existence of associativity pairs for non-associativity. For necessity, if any associator is non-zero then since $a=bcd$ then the associator must be ±$\pm 2a$. If only one associator type is zero and the other two are non-zero then from (10), the triple associator becomes
$T=\pm 2a\pm 2a+0=\pm 2a\text{,}$
which is again a contradiction and proves the conjecture. This theorem says nothing about all three associator types being non-zero and the triple associator is satisfied in this case, as well as satisfying the paired type associativity. ∎
Definition.
Non-associativity types are defined, using $\star$  to mean non-zero, as
Associativity Non-Associativity
Type 1 Type 2 Type 3
$\star$ 0 0 Type A
0 $\star$ 0 Type B
0 0 $\star$ Type C
$\star$ $\star$ $\star$ Type X
Definition.
Triad $(b,c,d)$ generates a Malcev algebra if it satisfies Mal’cev’s identity [10], which is immediately simplified for ultracomplex numbers,
$\begin{split}bc(bd)-bcdb-cdbb-dbbc=0,\\ -bc(db)-bcdb+cd+dc=0\text{ or}\\ bcdb+bc(db)=0\text{.}\end{split}$ (11)
The simplified identity means $[bc,d,b]\neq 0$ which implies A or X non-associativity if $bc$ is interpreted as a new $c$. Malcev algebras define non-associative Lie algebras and this is analysed later where B and C non-associativity types are also found. For quaternion-like triads, $d=bc$ can be substituted into (11),
$bcdb+bc(db)=bc(bc)b+bc(bcb)=-b+bcc=-2b$
and similarly, cycles $d=b$ or $d=c$ are also non-zero. So associative algebras are not Malcev algebras.
Theorem 6 (Moufang associativity theorem).
A Moufang loop [11] is a non-associative algebraic structure that satisfies three identities for non-associative triad $(b,c,d)$. Calling these types 1 to 3, then each type exhibits the non-associativity shown in the following table.
Type Moufang non-associativity Equivalence
1 $d(b(dc))-dbdc=0$ B or X
2 $b(d(cd))-bdcd=0$ C or X
3 $db(cd)-d(bc)d=0$ B or X if Malcev,
A or C otherwise
Proof.
Type 1 and 2 equations, multiplied by $d$ and assuming $d^{2}=-1$, become respectively
$d(b(dc))d-dbdcd$ $=-b(dc)dd+bddcd=dcbdd-bcd=-dcb-bcd=-bcd-b(cd)=0\text{ and}$
$b(d(cd))d-bdcdd$ $=-d(cd)bd+dbcdd=-cddbd+dbcdd=cbd-dbc=cdb+c(db)=0\text{.}$
Substituting the first equation into Type 2 associativity is $[b,c,d]=bcd-b(cd)=2bcd$, which is non-zero. Substituting the second into Type 3 associativity is $[c,b,d]=cbd-c(bd)=2cbd$, which is also non-zero. Hence Moufang loop non-associativity of Type 1 and 2 corresponds to B-non-associativity and C-non-associativity, respectively, or X-non-associativity in both cases.
Multiplying Moufang type 3 by $db$ gives,
$db(cd)(db)-d(bc)d(db)=-cd(db)(db)-bc(db)=0\text{,}$
so $cd=bc(db)$. Applying (11) so that $cd=-bcdb$ and substituting into Type 2 associativity gives,
$[b,c,d]=bcd-b(cd)=bcd+b(bcdb)=bcd-bcdbb=bcd+bcd=2bcd\neq 0\text{.}$
This implies type B or X non-associativity. If Mal’cev’s identity is not satisfied then (11) takes the opposite sign, $[bc,d,b]=0$, and Type 2 associativity is then zero so that A or C non-associativity holds. This completes the proof and it is noted that it is possible to use combinations of the Moufang types along with Mal’cev’s identity to distinguish all four types of non-associativity A, B, C and X. ∎
The associative triads have pairs $(b,c)$, $(b,bc)$ and $(c,bc)$, which form a 3-cycle of pairs. This applies to all pairs and can be extended to triads to provide another level of structure to the ultronions.
Table 2: Triad Structure
Label Associative Non-Cycles 3-Triad Cycles Total
${\mathbb{C}}$ 0 0 0 0
${\mathbb{H}}$ 1 0 0 1
${\mathbb{O}}$ 7 4 24 35
${\mathbb{U}_{1}}$ 35 60 360 455
${\mathbb{U}_{2}}$ 155 620 3,720 4,495
${\mathbb{U}_{3}}$ 651 5,580 33,480 39,711
${\mathbb{U}_{4}}$ 2,667 47,244 283,464 333,375
${\mathbb{U}_{5}}$ 10,795 388,620 2,331,720 2,731,135
${\mathbb{U}_{6}}$ 43,435 3,152,140 18,912,840 22,108,415
${\mathbb{U}_{7}}$ 174,251 25,390,860 152,345,160 177,910,271
Definition.
Naturally ordered basis pairs form 3-cycles, $(b,c)$, $(b,bc)$ and $(c,bc)$, which cover all ordered pairs. This is because there are $(2^{n}-1)(2^{n}-2)/2$ combinations of pairs which have factors one and two numbers less than a power of $2$ so one must have a factor of $3$. These pairs can be ordered as $b<c<bc$ and the pairs of the ordered 3-cycles are called the first, second and third cycle, respectively. A 3-triad cycle extends this to the triads $(b,c,d),(b,bc,d),(c,bc,d)$ where $b<c<bc<d$. These are not distinct since triads with $d<bc$ have second and third cycles already included elsewhere. Such cases are called non-cycles and associative triads, $d=bc$, are also degenerate.
Table 2 enumerates all triads for Cayley-Dickson algebras for up to $10$ generators. Non-associative triads are split into non-cycles and cycles where the non-cycle ones have $d<bc$. The summation of these triad enumerations is provided in the last column of Table 2 that matches $\binom{N}{3}$, where $N=2^{n}-1$.
For quaternions, ${\mathbb{H}}$, the associative triad in Table 2 is $(o_{1},o_{2},o_{12})$. For octonions, ${\mathbb{O}}$, the $7$ associative triads have been provided by (LABEL:eqn:qa7). The $24$ 3-triad cycles are derived from eight triads starting with the following first cycles,
$(o_{1},o_{2},o_{3})$, $(o_{1},o_{2},o_{13})$, $(o_{1},o_{2},o_{23})$, $(o_{1},o_{2},o_{123})$,
$(o_{1},o_{3},o_{23})$, $(o_{1},o_{3},o_{123})$, $(o_{2},o_{3},o_{123})$, $(o_{12},o_{13},o_{123})$.
All these octonion triads have X-non-associativity and it will soon be proved that the cycles must also be of type X. Remember $[o_{1},o_{2},o_{3}]=2o_{123}$ and multiplying terms by distinct basis elements does not change the non-associativity type. To achieve this we need another generator, $o_{4}$, for example, so that for $b=o_{1}$, $c=o_{2}$ and $d=o_{34}$ then Type 2 is associative, $[o_{2},o_{3},o_{34}]=-o_{1234}+o_{1234}=0$ and Type 1 is non-associative, $[o_{1},o_{34},o_{2}]=-o_{1234}-o_{1234}\neq 0$.
Table 3 provides the first A and C triads from ${\mathbb{U}_{1}}$ and the first B triad from ${\mathbb{U}_{2}}$ as examples. For X non-associativity, the first ${\mathbb{U}_{1}}$ X case not included in ${\mathbb{O}}$ is selected but note that any octonion non-associative triad could have been selected with generator $o_{3}$ replaced with $o_{4}$.
Table 3: Non-associativity to Associativity Examples
Type Type 1 Type 2 Type 3
A $[o_{1},o_{34},o_{2}]=-2o_{1234}$ $[o_{1},o_{2},o_{34}]=0$ $[o_{2},o_{1},o_{34}]=0$
B $[o_{1},o_{345},o_{24}]=0$ $[o_{1},o_{24},o_{345}]=-2o_{1235}$ $[o_{24},o_{1},o_{345}]=0$
C $[o_{1},o_{34},o_{24}]=0$ $[o_{1},o_{24},o_{34}]=0$ $[o_{24},o_{1},o_{34}]=-o_{123}$
X $[o_{1},o_{4},o_{2}]=-2o_{124}$ $[o_{1},o_{2},o_{4}]=2o_{124}$ $[o_{2},o_{1},o_{4}]=-o_{124}$
The enumeration of 3-triad cycles in Table 2 can be divided by three and categorised as belonging to one of eight 3-triad cycles and the enumeration of non-cycles is separated into the four single triads rows shown in Table 4. This table provides the 3-triad cycles structure for Cayley-Dickson algebras up to 10 generators and will be used to categorise zero divisors in the next section. As can be seen the structures up to ${\mathbb{U}_{4}}$ are distinct and after this the structure is the same so only ${\mathbb{U}_{1}}$, ${\mathbb{U}_{2}}$, ${\mathbb{U}_{3}}$ and ${\mathbb{U}_{4}}$ provide algebras with unique structure for ultracomplex numbers. The following theorem validates this table.
Theorem 7 (Cycles theorem).
Of the 64 possible non-associativity 3-triad cycles only silos AAA, ACC, XBB, BBA, BXC, CAB, CCX and XXX exist.
Table 4: Non-associative Structure
Structure ${\mathbb{O}}$ ${\mathbb{U}_{1}}$ ${\mathbb{U}_{2}}$ ${\mathbb{U}_{3}}$ ${\mathbb{U}_{4}}$ ${\mathbb{U}_{5}}$ ${\mathbb{U}_{6}}$ ${\mathbb{U}_{7}}$
AAA 0 28 252 1,988 14,868 111,300 849,492 6,605,284
BBA 0 0 168 1,848 15,120 115,080 872,928 6,721,848
ACC 0 0 84 1,344 12,852 105,672 834,876 6,569,136
XBB 0 0 0 672 9,072 87,360 752,976 6,219,360
BXC 0 0 0 168 4,536 60,648 621,432 5,628,168
CAB 0 0 0 0 3,024 31,744 577,584 5,431,104
CCX 0 28 336 2,660 18,648 129,612 931,392 6,955,060
XXX 8 64 400 2,480 16,368 115,824 863,600 6,651,760
A 0 0 84 1,008 9,828 87,864 747,180 6,175,008
B 0 0 0 504 7,560 78,456 709,128 6,022,296
C 0 28 252 1,988 14,868 111,300 849,492 6,605,284
X 4 32 284 2,080 14,988 111,000 856,340 6,588,272
Proof.
Starting with A-non-associativity for the first cycle then showing that the second and third cycles have Type 2 associativity shows that these are either A or C-non-associative. Here we use Type 2 associativity $[b,c,d]=0$ and Type 3 associativity $[c,b,d]=0$ and apply to the second cycle equivalent,
$[b,bc,d]$ $=b(bc)d-b(bcd)=-bcbd+bcdb=cbbd+b(cd)b=-cd+cd=0\text{ and}$
$[c,bc,d]$ $=c(bc)d-c(bcd)=-bccd+bcdc=bd-c(bd)c=bd-bd=0\text{.}$ (13)
Hence the result is either A or C and AA means third cycle A and AC means third cycle C needs to be shown. Using Type 3 associativity, $[c,b,d]=0$, then third cycle Type 2 associativity is
$[c,bc,d]=c(bc)d-c(bcd)=-bccd-cbdc=bd-c(bd)c=0\text{,}$
meaning either AAA or AAC. Now assume third cycle Type 3 associativity, then
$[bc,c,d]=bccd-bc(cd)=0\implies db=-bc(cd)$
Substitute this into second Type 3 associativity,
$[bc,b,d]=bcbd-bc(db)=-cbbd+bc(bc(cd))=cd+cd(bc)(bc)=0\text{,}$
which means the third cycle for AA is also A. To the contrary, if the second cycle is not Type 3 associative then neither is the third so it must be ACC.
For the BBA and BXC silos we need to show that the second cycle has non-zero Type 2 associator and third cycle Type 2 associativity. Starting with the Type 2 associator for the second and third cycles and using Type 3 associativity, $[c,b,d]=0$, and Type 1 associativity, $[b,d,c]=0$, then
$[b,bc,d]$ $=b(bc)d-b(bcd)=-bcbd+bcdb=cbbd-c(bd)b=dc+b(dc)b=2dc\neq 0\text{ and}$
$[c,bc,d]$ $=c(bc)d-c(bcd)=-bccd+bcdc=bd-c(bd)c=bd-bd=0\text{.}$
So the second cycle is either B or X-non-associative and the third A or C. If the second cycle is B then the third can be expanded using Type 3 and Type 1 associativity. Considering third cycle Type 1 associativity and using Type 3 associativity and assuming the third cycle Type 3 associativity from (13), then
$[c,d,bc]=cd(bc)-c(d(bc))=cd(bc)-bcdc=cd(bc)+c(bd)c=cd(bd)+bc(cd)=cd(bd-bc)\neq 0\text{.}$
Hence BB has A as the third cycle and therefore BXC is the other 3-cycle.
For CAB and CCX silos, consider the second and third Type 2 associativity. Using Type 2 associativity, $[b,c,d]=0$, and Type 1 associativity, $[b,d,c]=0$, then
$[b,bc,d]$ $=b(bc)d-b(bcd)=-bcbd+bcdb=cbbd+b(cd)b=-cd+cd=0\text{ and}$
$[c,bc,d]$ $=c(bc)d-c(bcd)=-bccd+bcdc=bd+b(cd)c=bd-bdcc=2bd\neq 0\text{.}$
Hence the second cycle is either A or C and the third either B or X. This is separated by showing that the second and third associator are either both associative or both non-zero. Substituting second Type 3 associativity, $[bc,b,d]=0$ as $cd=bc(db)$ into the third gives,
$[bc,c,d]=bccd-bc(cd)=-bd-bc(bc(db))=-bd-db(bc)(bc)=0\text{.}$
Hence C in the first cycle leads to CAB or CCX silos.
The silos that start with X non-associativity are proved using Moufang’s non-associativity. From Theorem 6, for X-non-associativity the Moufang types are 1 and 2,
$d(b(dc))-dbdc=dcbd+bddc=0$ $\quad\Rightarrow\quad bc=dcbd\text{ and}$ (14)
$b(d(cd))-bdcd=cddb-bdcd=0$ $\quad\Rightarrow\quad bc=bdcd\text{.}$ (15)
So the second cycle is either B or X. Type 3 and 1 for the second cycle using (15) for Type 3 is
$[bc,b,d]$ $=bcbd-bc(bd)=cd-bc(bd)\text{,}$ (16)
$[b,d,bc]$ $=bd(bc)-b(d(bc))=bd(bc)-bcdb=bd(bc)-dcbddb=bd(bc)-dc=[bc,b,d]\text{.}$
Type 2 and 3 for the third cycle with (14) applied to Type 2 is
$[c,bc,d]$ $=c(bc)d-c(bcd)=bd+bcdc=bd+bdcddc=2bd\neq 0\text{,}$
$[bc,c,d]$ $=bccd-bc(cd)=-bd-bc(cd)\text{.}$ (17)
Hence, the third cycle is B or X, since Type 2 is non-zero, assuming the first cycle is X or B, from (14). If Type 3 second cycle is zero, such that $cd=bc(bd)$ from (16), then (17) becomes
$-bd-bc(cd)=-bd+bc(bd)(bc)=-bd-bd(bc)(bc)=0\text{.}$
So again, the third cycle is B if the second is B, alternatively, both are X. Since the first cycle was assumed to be X we have 3-triads cycles XXX or XBB. ∎
Half of the 3-triad cycle silos satisfy the identity for Malcev algebra, (11), which can now be recognised as second and third cycles from the Cycle theorem. Due to the cyclic rules the AAA and XXX silos satisfy (11) for all cycles so each 3-triads cycle provides a Malcev algebra of 3 triads. This is applicable to the eight XXX cycles of ${\mathbb{O}}$ indicated in Table 4 and shown in (12). The four non-cycle X triads are $(o_{2},o_{3},o_{13})$, $(o_{2},o_{13},o_{23})$, $(o_{12},o_{3},o_{13})$ and $(o_{12},o_{3},o_{23})$ and these also satisfy Mal’cev’s identity since their second and third cycle parts overlap with other X triads. This completes the 28 Malcev triads in ${\mathbb{O}}$ that provide the tangent space of the 7-sphere since the 7 associative ${\mathbb{O}}$ triads do not satisfy the Malcev identity. There are four remaining silos with partial Malcev properties. The ACC and XBB satisfy (11) for the second and third cycles but not the first. The other two cases are BAC and CAB where the first cycle are Malcev algebra triads but the second and third are not. All these cases are derived from (11) which implies that only cycles that include A and X cases will satisfy this identity.
The X non-associative triad is the only completely symmetric non-associative type and this is the characteristic of octonions. This identifies the non-associative triads of ${\mathbb{O}}$ in the XXX silo and the ${\mathbb{O}}$ parts of ${\mathbb{U}_{1}}$ in Table 4. Each triad generates 7 distinct blades giving $\binom{7}{3}=28$ distinct triads and the members of 3-triad cycles are already included in this set so only the first cycle needs to be considered. ${\mathbb{O}}$ have $8\times 3=24$ triads in the XXX silo and 4 in the non-cycles. ${\mathbb{U}_{1}}$ has eight times these numbers and these can be identified with 8 triads that generate the composed ${\mathbb{O}}$ then seven copies of ${\mathbb{O}}$,
$\begin{split}&(o_{1},o_{2},o_{3}),(o_{1},o_{2},o_{4}),(o_{1},o_{3},o_{4}),(o_{1},o_{23},o_{4}),\\ &(o_{2},o_{3},o_{4}),(o_{2},o_{13},o_{4}),(o_{12},o_{3},o_{4}),(o_{12},o_{13},o_{4})\text{.}\end{split}$
The remaining AAA and CCX silos have 56 3-triad cycles giving seven copies of another algebra with triads
$(o_{1},o_{2},o_{34}),(o_{1},o_{3},o_{24}),(o_{1},o_{23},o_{24}),(o_{2},o_{3},o_{14}),(o_{2},o_{13},o_{14}),(o_{12},o_{3},o_{14}),(o_{12},o_{13},o_{14})\text{.}$
This is isomorphic to an algebra documented in [12] which is part of a series of six algebras that break the symmetry of octonions whereby subalgebras of exceptional Lie algebra G2 provide automorphisms for these algebras. These have properties similar to the split octonions seen later, so were called the pseudo-octonions, designated ${\mathbb{P}_{k}}$ where $k$ is the number of non-associative triads $k=4,8,10,12,14,16$. The algebra above is ${\mathbb{P}_{4}}$ and only two other pseudo-octonion algebras, ${\mathbb{P}_{12}}$ and ${\mathbb{P}_{14}}$, are subalgebras of ultracomplex numbers. The number of occurrences of each subalgebra for the same algebras as shown in Table 4 is provided in Table 5.
Table 5: Ultracomplex Subalgebras
Structure ${\mathbb{O}}$ ${\mathbb{U}_{1}}$ ${\mathbb{U}_{2}}$ ${\mathbb{U}_{3}}$ ${\mathbb{U}_{4}}$ ${\mathbb{U}_{5}}$ ${\mathbb{U}_{6}}$ ${\mathbb{U}_{7}}$
${\mathbb{H}}$ 7 35 155 651 2,667 10,795 43,435 174,251
${\mathbb{O}}$ 1 8 50 310 2,046 14,478 107,950 831,470
${\mathbb{P}_{4}}$ 0 7 63 413 2,583 16,905 118,251 873,901
${\mathbb{P}_{12}}$ 0 0 42 504 4,158 30,996 229,194 1,729,728
${\mathbb{P}_{14}}$ 0 0 0 168 3,024 34,776 332,640 2,912,616
Total ${\mathbb{P}_{k}}$ 0 7 105 1085 9,765 82,677 680,085 5,516,245
Theorem 8 (Ultronion subalgebras).
Defining $S_{m}$ and $O_{m}$ as the number of total ${\mathbb{P}_{k}}$ and ${\mathbb{O}}$ copies in ${\mathbb{U}_{m}}$ where the number of generators is $m+3$ then $S_{m+1}=7\times(O_{m}+S_{m})$. For convenience, $m=0$ case is taken to be ${\mathbb{U}_{0}}={\mathbb{O}}$, so $m\geq 0$.
Proof.
Table 5 shows these numbers for cases $m\in{\mathbb{N}}_{0}^{7}$. The number of blades or pure basis elements is $N_{m}=2^{m+3}-1$ so that the number of ${\mathbb{H}}$ in ${\mathbb{U}_{m}}$ is given by the number of paired blades, $\binom{N_{m}}{2}$, as seen in the table. The total of the number of triads in ${\mathbb{U}_{m}}$ is $\binom{N_{m}}{3}$ which is observed from Table 5 to be
$\binom{N_{m}}{3}=O_{m}+S_{m}+\binom{N_{m}}{2}\text{.}$
It is also observed that each ${\mathbb{P}_{k}}$ is divisible by 7, hence $S_{m}$ is divisible by 7, and that $S_{m}=(S_{m-1}+O_{m-1})\times 7$. Hence
$\frac{S_{m}}{S_{m-1}}=\frac{\binom{N_{m}}{3}-\binom{N_{m}}{2}}{\binom{N_{m-1}}{3}-\binom{N_{m-1}}{2}}\text{,}$
so the ratio of increase of the pseudo-octonion algebras equates to the rate of increase of the non-associate triads. By induction the series proceeds past ${\mathbb{U}_{7}}$ with the same ratio. Thus, the increase from ${\mathbb{U}_{m}}$ to ${\mathbb{U}_{m+1}}$ is $S_{m+1}=7\times(O_{m}+S_{m})$ and the octonions fill in the difference, $O_{n+1}=\frac{S_{n+2}}{7}-S_{n+1}$. ∎
The silo sources for each subalgebra are provided in Table 6 along with the non-associative type counts. Each algebra has 8 3-triad cycles from the silos totalling 24 triads and the remaining 4 triads are always found in the non-cycles, so these are not identified but can be discovered by subtracting the identification column from the silos. It is interesting to note that all subalgebras have four X non-associativity types so a selection of a triad of type X does not guarantee an algebra isomorphic to ${\mathbb{O}}$ will be generated.
Table 6: Subalgebra Identification
Subalgebra Identification Silo counts
${\mathbb{O}}$ 28X 8XXX
${\mathbb{P}_{4}}$ 12A, 12C, 4X 4AAA and 4CCX
${\mathbb{P}_{12}}$ 8A, 8B, 8C, 4X 4BBA, 2ACC, 2CCX or
2AAA, 2BBA, 2BXC, 2CCX
${\mathbb{P}_{14}}$ 7A, 10B, 7C, 4X AAA, 3ACC, 4XBB or
AAA, XBB, 3BXC, 3CAB or
BBA, 2ACC, 2XBB, BXC, 2CAB or
BBA, ACC, XBB, 2BXC, 3CAB
Theorem 9 (Projection Theorem).
Any triad from an ultracomplex algebra will generate an algebra isomorphic or anti-isomorphic to ${\mathbb{H}}$, ${\mathbb{O}}$, ${\mathbb{P}_{4}}$, ${\mathbb{P}_{12}}$ or ${\mathbb{P}_{14}}$. Furthermore, triads from ${\mathbb{U}_{1}}$ can project (or anti-project) to ${\mathbb{H}}$, ${\mathbb{O}}$ or ${\mathbb{P}_{4}}$, from ${\mathbb{U}_{2}}$ to a ${\mathbb{U}_{1}}$ subalgebra or ${\mathbb{P}_{12}}$ and all higher levels can project to ${\mathbb{U}_{2}}$ subalgebras or ${\mathbb{P}_{14}}$. This leads to the diagram of injections shown in Figure 1.
Figure 1: Injections of Cayley Dickson algebras
Proof.
The injections of the normed composition algebras, ${\mathbb{C}}$, ${\mathbb{H}}$, ${\mathbb{O}}$, are well defined by doubling, up to ${\mathbb{O}}$, which has 7 copies of ${\mathbb{H}}$ and a new non-associative part given by 28 triads. These correspond to the faces of the 6-simplex and the non-face triads, respectively, as discussed in [12]. The associative and non-associative parts can be isolated, with associative triads being quaternion-like by Theorem 3, and the non-associative injection into the ultronions analysed separately. The number of distinct ${\mathbb{O}}+{\mathbb{P}_{k}}$ subalgebra copies for each of the listed ultronions matches the number of total non-associative triads divided by 28, as shown in Table 14, hence these subalgebra copies provide complete coverage of each ultracomplex algreba. A member of the XXX silo, $(o_{1},o_{2},o_{3})$, generates ${\mathbb{O}}$ and the first ordered member of CCX, $(o_{1},o_{24},o_{34})$, generates ${\mathbb{P}_{4}}$. Table 4 shows that these are the only silos available for ${\mathbb{U}_{1}}$. Then ${\mathbb{U}_{2}}$ contains these silos and BBA with the first ordered member, $(o_{1},o_{24},o_{345})$, generating ${\mathbb{P}_{12}}$. After this ${\mathbb{U}_{3}}$ contains silo XBB with first ordered member $(o_{14},o_{25},o_{36})$, which generates ${\mathbb{P}_{14}}$. Higher level ultracomplex algebras contain CAB, which can only generate ${\mathbb{P}_{14}}$, for example the first member that first appears in ${\mathbb{U}_{7}}$, $(o_{14},o_{256},o_{347})$. ∎
As shown later, each of the pseudo-octonion algebras have 12 zero divisors making them power-associative so this property flows through to the ultracomplex algebras. Since there are always seven copies of any of the pseudo-octonion algebras then zero divisors appear in multiples of 84.
4 Zero Divisors
Definition.
A Zero divisor is a term whereby another term exists such that both are non-zero and the product is zero. Zero divisors can not be single blades so a zero divisor pair is defined, $(a+b)(c+d)=0$, for $a,b,c,d$ all distinct pure scaled basis elements and $a^{2}=b^{2}$ and $c^{2}=d^{2}$.
Theorem 10 (Zero divisors theorem).
The parallel terms and cross terms of the zero divisors equation separate into
$ac+bd=0\text{ and }ad+bc=0\text{,}$ (18)
with the second equation being an identity thus reducing the definition to just $ac=-bd$. Further, this is equivalent to Type 3 associativity and means the number of zero divisors at level $n$, $Z_{n}$, is
$Z_{n+1}=\left(\binom{2^{n}-1}{3}-\binom{2^{n}-1}{2}\right)\times 84\text{.}$
Proof.
Uniqueness of $a,b,c$ and $d$ ensures $ac$ does not have the same blade as either $ad$ or $bc$, from the symmetric difference of Lemma 1, so the two equations (18) have distinct scaled blades and are separable. Since $b,c,d$ are distinct blades, they anticommute, and since $a$ is pure then $a\approx bcd$ then $a$ is non-scalar so that $b,c,d$ are non-associative, $d\neq bc$. Hence triple products also anticommute and the equations can be re-arranged as
$a=-(bd)c^{-1}=c(bd)/c^{2}$ and $a=-(bc)d^{-1}=cbd/d^{2}\text{ and }$ and (19)
[…]
which implies $c^{2}=d^{2}$ and $a^{2}=b^{2}$. Substituting (19) and (20) shows Type 3 associativity
$\begin{split}[b,c,a]&=bca-b(ca)=da^{2}-db^{2}=0\text{ and}\\ [c,b,d]&=cbd-c(bd)=ad^{2}-ac^{2}=0\text{.}\end{split}$
Note that the reversed product $(c+d)(a+b)$ generates $-ac-bd=0$ so provide the same zero divisor pair. Finally, substituting for $a$, $ad=cbdd/d^{2}=-bc$ so that $ad+bc=0$ is an identity.
Hence zero divisor pairs are separated into two types for A and B non-associative triads that are both Type 3 associative. Octonions can not have zero divisors because the non-associative triads are X non-associative. Swapping the last two elements of the associator in Table 1 swaps Type 1 associativity and Type 2 associativity and has no correspondence with Type 3 associativity other than changing the sign. Hence A and B zero divisors may overlap because swapping the terms within brackets produces the same zero divisor pair. Table 5 shows 12 A non associative types for ${\mathbb{P}_{4}}$, which corresponds to 12 zero divisor pairs. The other pseudo-octonion algebras have mixures of A and B types adding up to more than 12 but since all ordered triples are included in the list of 28 triads then some A and B types overlap as zero divisors to give 12 distinct zero divisor pairs. Theorem 8 defines the number of pseudo-octonion algebras at each level as $S_{m+1}=7\times(O_{m}+S_{m})$ where $O_{m}$ and $S_{m}$ are the number of octonion and pseudo-octonion algebras at level $n=m+3$, respectively. This leads to $Z_{n+1}=(O_{n}+S_{n})\times 7\times 12$. ∎
Definition (Modes).
Calling (18) the prime zero divisor pair then $(-d+b)(c+a)$, is called the dual of the prime. The terminology dual is appropriate for zero divisors. Also $(a^{\prime}+b)(c+\mathop{\rm abs}\nolimits(db))$, with $a^{\prime}=bc\mathop{\rm abs}\nolimits(db)$, is called the extended zero divisor pair, and $(-\mathop{\rm abs}\nolimits(db)+b)(c+a^{\prime})$, is called both extended and dual. These four cases are called modes and whether they are zero divisors determines another level of structure for the ultronions.
Theorem 11 (Modes theorem).
All dual mode transformations swap non-associativity types between pairs of A, pairs of X and between B and C. All non-cycle extended and extended dual transformations may act like the dual transformation or as shown in Figure 2. Transformations of 3-triad cycles are as described in the following table.
Type Related Types
AAA All modes transform like the dual mode to other A types
BBA All modes transform as shown in Figures (a) and (b)
ACC All modes transform as shown in Figures (a) and (c)
XBB All modes transform as shown in Figures (a) and (c)
BXC All modes transform as shown in Figures (a) and (c)
CAB All modes transform as shown in Figures (a) and (c)
CCX All modes transform as shown in Figures (a) and (b)
XXX All modes transform like the dual mode to other X types
Figure 2: Special mode transformations
Proof.
The dual mode replaces $d$ with $a$ in the triad $(b,c,d)$ so substituting this into Table 1 we find,
Type 1 associativity [$\text{Type\leavevmode\nobreak\ 1 associativity }[b,d,c]$ $\rightarrow[b,a,c]\text{ which is Type\leavevmode\nobreak\ 1 associativity,}$,
[…]
Together these mean under the dual transformation A $\rightarrow$ A, X $\rightarrow$ X and B $\leftrightarrow$ C and this applies to all the triad cycles and to dual extended mode as well. Hence, only the extended mode, where $d\rightarrow\mathop{\rm abs}\nolimits(db)$, for all cycles needs to be considered.
For easier traceability, the terminology ${\rm P}p_{q}$ is introduced to identify a prime Type $p$ associativity equation for the $q^{\rm th}$ cycle, where $p,q\in{\mathbb{N}}_{1}^{3}$ and ${\rm E}p_{q}$ for the equivalent types and cycles for the extended modes. Then the Type 1 and Type 3 associativity equations for the first and second cycles are
[…]
The extended mode associativity equations that are required are Types 1 and 3 for the first and second cycles and Types 2 and 3 for the third cycle,
[…]
Table 7: Extended Mode Transformations
Prime Mode Extended Mode
[…]
AAA $\star$ 0 $\star$ 0 $\star$ 0 $\star$ 0 0 0 AAA
BBA 0 0 0 0 0 0 0 0 0 0 BBA
ACC $\star$ 0 0 $\star$ $\star$ $\star$ 0 0 $\star$ 0 XBB
XBB $\star$ $\star$ 0 0 $\star$ 0 0 $\star$ 0 $\star$ ACC
BXC 0 0 $\star$ $\star$ 0 $\star$ $\star$ 0 $\star$ 0 CAB
CAB 0 $\star$ $\star$ 0 0 0 $\star$ $\star$ 0 $\star$ BXC
CCX 0 $\star$ 0 $\star$ 0 $\star$ 0 $\star$ $\star$ $\star$ CCX
XXX $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ XXX
It is observed that the non-cycle transformations occur in blocks according to the cycle $(b,c)$ with $d<bc$. The A and X transformations for non-cycles act like the dual whereas overlapping cycle cases may swap types but keep the same number of each type for each cycle. ∎
The definition of zero divisor pairs required that $a^{2}=b^{2}$ and $c^{2}=d^{2}$. For the extended equation to hold these become $a^{\prime 2}=b^{2}$ and $c^{2}=(db)^{2}$, which implies $b^{2}=c^{2}=-1$. Hence for all subsequent work we assume $a^{2}=b^{2}=c^{2}=d^{2}=\pm 1$ and take $b$, $c$ and $d$ to be positive blades.
All zero divisor pairs in Table 4 are derived only from the silos that contain A and B non-associative types. The mode transformations for the AAA silo all act as dual transformation, which by Figure 2(a) provides other A types. Hence the AAA silo can be reduced by factors of 3 cycles and 4 modes in all ultracomplex algebras. The BBA silo can also be reduced by 3 cycles but the modes are more complicated. Only the A dual mode and B extended mode generate zero divisors so modes provide a reduction by two only. It is observed that for the ultronions shown in Table 4, that for ${\mathbb{U}_{1}}$ only the AAA silo can provide distinct zero divisor pairs, numbering $28\times 3=84$.
Cawagas [5] labels the pure sedenions as $e_{i}$, for $i\in{\mathbb{N}}_{1}^{16}$. The graded single, pure elements, called the basis of ${\mathbb{S}}\equiv{\mathbb{U}_{1}}$, have a natural order generated by the Cayley-Dickson process, which match these in order as
$(e_{1},e_{2},\dots e_{15})=(o_{1},o_{2},o_{12},o_{3},o_{13},o_{23},o_{123},o_{4},o_{14},o_{24},o_{124},o_{34},o_{134},o_{234},o_{1234})\text{.}$ (21)
The first three elements are recognised as the pure quaternions and the first seven match the octonions [2]. Cawagas derives the $84$ zero divisor pairs shown in [5] and these are provided in graded form in Table 8 as $b$ and $c$ columns followed by prime, dual, extended and its dual triads as $d$, $a$, $db$ and $ab$ columns, respectively.
Sedenions are power-associative which means the associator with repeated terms is always zero, $[a,a,a]=0$. This is obvious for an algebra where the single terms square to a scalar and other than the zero divisor pairs all terms have inverses. Octonions are alternate associative which means swapping terms of the associator changes the sign or $[a,a,b]=[a,b,a]=0$ for all $a,b\in{\mathbb{O}}$. But sedenions are not alternate associative and demonstrate asymmetric associativity. For example, the A type from Table 3 is
$\left[o_{1},o_{2},o_{34}\right]$ $=-o_{12}o_{34}-o_{1}o_{234}=0\text{ and }$
$\left[o_{1},o_{34},o_{2}\right]$ $=o_{134}o_{2}+o_{1}o_{234}=-2o_{1234}\text{.}$
Table 8: ${\mathbb{U}_{1}}$ Zero Divisors
b c d a db ab
1 $o_{1}$ $o_{2}$ $o_{34}$ $-o_{1234}$ $o_{134}$ $o_{234}$
2 $o_{1}$ $o_{12}$ $o_{34}$ $o_{234}$ $o_{134}$ $o_{1234}$
3 $o_{1}$ $o_{3}$ $o_{24}$ $o_{1234}$ $o_{124}$ $-o_{234}$
4 $o_{1}$ $o_{13}$ $o_{24}$ $-o_{234}$ $o_{124}$ $-o_{1234}$
5 $o_{1}$ $o_{23}$ $o_{24}$ $o_{134}$ $o_{124}$ $o_{34}$
6 $o_{1}$ $o_{123}$ $o_{24}$ $-o_{34}$ $o_{124}$ $o_{134}$
7 $o_{2}$ $o_{12}$ $o_{34}$ $-o_{134}$ $o_{234}$ $o_{1234}$
8 $o_{2}$ $o_{3}$ $o_{14}$ $-o_{1234}$ $-o_{124}$ $-o_{134}$
9 $o_{2}$ $o_{13}$ $o_{14}$ $o_{234}$ $-o_{124}$ $o_{34}$
10 $o_{2}$ $o_{23}$ $o_{14}$ $-o_{134}$ $-o_{124}$ $o_{1234}$
11 $o_{2}$ $o_{123}$ $o_{14}$ $o_{34}$ $-o_{124}$ $-o_{234}$
12 $o_{12}$ $o_{3}$ $o_{14}$ $o_{234}$ $o_{24}$ $-o_{134}$
13 $o_{12}$ $o_{13}$ $o_{14}$ $o_{1234}$ $o_{24}$ $o_{34}$
14 $o_{12}$ $o_{23}$ $o_{14}$ $-o_{34}$ $o_{24}$ $o_{1234}$
15 $o_{12}$ $o_{123}$ $o_{14}$ $-o_{134}$ $o_{24}$ $-o_{234}$
16 $o_{3}$ $o_{13}$ $o_{24}$ $-o_{124}$ $-o_{234}$ $-o_{1234}$
17 $o_{3}$ $o_{23}$ $o_{14}$ $o_{124}$ $-o_{134}$ $o_{1234}$
18 $o_{3}$ $o_{123}$ $o_{14}$ $-o_{24}$ $-o_{134}$ $-o_{234}$
19 $o_{13}$ $o_{23}$ $o_{14}$ $o_{24}$ $o_{34}$ $o_{1234}$
20 $o_{13}$ $o_{123}$ $o_{14}$ $o_{124}$ $o_{34}$ $-o_{234}$
21 $o_{23}$ $o_{123}$ $o_{24}$ $o_{124}$ $o_{34}$ $o_{134}$
This example produces the zero divisor pair $(o_{1}-o_{1234})(o_{2}+o_{34})$ that provides an example of non-alternate associativity, $[a,a,b]\neq 0$ even though $[a,b,a]=0$,
$\left[(o_{1}-o_{1234}),(o_{1}-o_{1234}),(o_{2}+o_{34})\right]$ $=-2o_{2}-2o_{34}\text{ and }$
$\left[(o_{1}-o_{1234}),(o_{2}+o_{34}),(o_{1}-o_{1234})\right]$ $=0\text{.}$
This is because each term is invertible so that $(o_{1}-o_{1234})^{2}=-2$ but multiplying the zero divisor terms first gives zero. For ${\mathbb{U}_{1}}$, we found above that zero divisors need $d=o_{i4}$ or $d=o_{ij4}$ or $d=o_{1234}$, for $i,j\in{\mathbb{N}}_{1}^{3}$, but $b$ and $c$ do not contain $o_{4}$ in order to satisfy $[c,b,d]=0$. This means $a$ also contains the $o_{4}$ generator, satisfying the symmetry required for the dual and other modes.
Table 9: AAA Primary Zero Divisors in ${\mathbb{U}_{1}}$
bc d
$o_{1}o_{2}$ $o_{34}$
$o_{1}o_{3}$ $o_{24}$
$o_{1}o_{23}$ $o_{24}$
$o_{2}o_{3}$ $o_{14}$
$o_{2}o_{13}$ $o_{14}$
$o_{12}o_{3}$ $o_{14}$
$o_{12}o_{13}$ $o_{14}$
Sedenions contain $455$ ordered triples split into 35 associative and, from Table 4, $84$ A non-associative, $112$ C non-associative and $224$ X or completely non-associative. Cawagas’s 84 zero divisor pairs are summarised in AAA 3-triad cycle form in Table 9 which shows the primes only. Applying the Mode and Cycles Theorems gives twelve zero divisor pairs, applicable for the AAA silo, over the seven 3-triad cycles generating all $84$ zero divisor pairs in Table 8.
Table 10: AAA Primary Zero Divisors in ${\mathbb{U}_{2}}$ based on ${\mathbb{O}}$
bc cycle d d1 d2
$o_{1}o_{2}$ $o_{1}o_{12}$ $o_{2}o_{12}$ $o_{34}$ $o_{35}$ $o_{45}$
$o_{1}o_{3}$ $o_{1}o_{13}$ $o_{3}o_{13}$ $o_{24}$ $o_{25}$ $o_{45}$
$o_{1}o_{23}$ $o_{1}o_{123}$ $o_{23}o_{123}$ $o_{24}$ $o_{25}$ $o_{45}$
$o_{2}o_{3}$ $o_{2}o_{23}$ $o_{3}o_{23}$ $o_{14}$ $o_{15}$ $o_{45}$
$o_{2}o_{13}$ $o_{2}o_{123}$ $o_{13}o_{123}$ $o_{14}$ $o_{15}$ $o_{45}$
$o_{12}o_{3}$ $o_{12}o_{123}$ $o_{3}o_{123}$ $o_{14}$ $o_{15}$ $o_{45}$
$o_{12}o_{13}$ $o_{12}o_{23}$ $o_{13}o_{23}$ $o_{14}$ $o_{15}$ $o_{45}$
The Cayley-Dickson algebra with five generators, ${\mathbb{U}_{2}}$, has $4495$ ordered triples split into $155$ associative, $1092$ B, $336$ A, $1092$ C and $1820$ X non-associative triads. Now ${\mathbb{U}_{2}}$ has $1260$ zero divisor pairs because unlike ${\mathbb{U}_{1}}$ with only AAA 3-triad cycles, Table 4 shows that BBA and ACC cases are involved. The AAA and BBA silos provide all the zero divisors so that the $84$ A types from the first cycle of the ACC silo and the $84$ non-cycle A types must overlap with the first two silos. So only the first two silos are analysed. What is not documented is that some of the BBA and AAA cases overlap, in that the same $b$, $c$ 3-cycle can be included in both silos for different $d$ elements. All these cases will be shown in different tables starting with ones based on octonion 3-cycles.
Table 11: AAA Primary Zero Divisors in ${\mathbb{U}_{2}}$ independent of ${\mathbb{O}}$
bc cycle d1 d2 d3
$o_{1}o_{4}$ $o_{1}o_{14}$ $o_{4}o_{14}$ $o_{25}$ $o_{35}$ $o_{235}$
$o_{2}o_{4}$ $o_{2}o_{24}$ $o_{4}o_{24}$ $o_{15}$ $o_{35}$ $o_{135}$
$o_{12}o_{4}$ $o_{12}o_{124}$ $o_{4}o_{124}$ $o_{15}$ $o_{35}$ $o_{135}$
$o_{3}o_{4}$ $o_{3}o_{34}$ $o_{4}o_{34}$ $o_{15}$ $o_{25}$ $o_{125}$
$o_{13}o_{4}$ $o_{13}o_{134}$ $o_{4}o_{134}$ $o_{15}$ $o_{25}$ $o_{125}$
$o_{23}o_{4}$ $o_{23}o_{234}$ $o_{4}o_{234}$ $o_{15}$ $o_{25}$ $o_{125}$
$o_{123}o_{4}$ $o_{123}o_{1234}$ $o_{4}o_{1234}$ $o_{15}$ $o_{25}$ $o_{125}$
Table 10 shows solutions derived from ${\mathbb{U}_{1}}$ which used the 4th generator in the $d$ term. It uses the Cycle and Mode Theorems to simplify the 84 zero divisor pairs of ${\mathbb{U}_{1}}$ to just 7 primaries and since they are AAA non-associative then all four modes can be derived from column $d$. The $bc$ 3-cycles cover all octonion pairs and as mentioned above only the $d$ term contains $o_{4}$. This is extended by ${\mathbb{U}_{2}}$ to include the 5th generator then both together, $o_{45}$, as shown in the $d_{1}$ and $d_{2}$ columns of the table. All of these triads are AAA non-associative with 3-cycle solutions thus giving three lots of $84$ or $252$ zero divisor pairs. The remaining AAA cases not based on octonion 3-cycles are provided in Table 11.
Table 12: Overlapping AAA and BBA Primary Zero Divisors in ${\mathbb{U}_{2}}$
bc cycle d d1 d2
$o_{1}o_{24}$ $o_{1}o_{124}$ $o_{24}o_{124}$ $o_{25}$ $o_{345}$ $o_{2345}$
$o_{1}o_{34}$ $o_{1}o_{134}$ $o_{34}o_{134}$ $o_{35}$ $o_{245}$ $o_{2345}$
$o_{1}o_{234}$ $o_{1}o_{1234}$ $o_{234}o_{1234}$ $o_{235}$ $o_{245}$ $o_{345}$
$o_{2}o_{14}$ $o_{2}o_{124}$ $o_{14}o_{124}$ $o_{15}$ $o_{345}$ $o_{1345}$
$o_{2}o_{34}$ $o_{2}o_{234}$ $o_{34}o_{234}$ $o_{35}$ $o_{145}$ $o_{1345}$
$o_{2}o_{134}$ $o_{2}o_{1234}$ $o_{134}o_{1234}$ $o_{135}$ $o_{145}$ $o_{345}$
$o_{12}o_{14}$ $o_{12}o_{24}$ $o_{14}o_{24}$ $o_{15}$ $o_{345}$ $o_{1345}$
$o_{12}o_{34}$ $o_{12}o_{1234}$ $o_{34}o_{1234}$ $o_{35}$ $o_{145}$ $o_{1345}$
$o_{12}o_{134}$ $o_{12}o_{234}$ $o_{134}o_{234}$ $o_{135}$ $o_{1245}$ $o_{345}$
$o_{3}o_{14}$ $o_{3}o_{134}$ $o_{14}o_{134}$ $o_{15}$ $o_{245}$ $o_{2345}$
$o_{3}o_{24}$ $o_{3}o_{234}$ $o_{24}o_{234}$ $o_{25}$ $o_{145}$ $o_{1245}$
$o_{3}o_{124}$ $o_{3}o_{1234}$ $o_{124}o_{1234}$ $o_{125}$ $o_{145}$ $o_{245}$
$o_{13}o_{14}$ $o_{13}o_{34}$ $o_{14}o_{34}$ $o_{15}$ $o_{245}$ $o_{1245}$
$o_{13}o_{24}$ $o_{13}o_{1234}$ $o_{24}o_{1234}$ $o_{25}$ $o_{145}$ $o_{1245}$
$o_{13}o_{124}$ $o_{13}o_{234}$ $o_{124}o_{234}$ $o_{125}$ $o_{145}$ $o_{245}$
$o_{23}o_{14}$ $o_{23}o_{1234}$ $o_{14}o_{1234}$ $o_{15}$ $o_{245}$ $o_{1245}$
$o_{23}o_{24}$ $o_{23}o_{34}$ $o_{24}o_{34}$ $o_{25}$ $o_{145}$ $o_{1345}$
$o_{23}o_{124}$ $o_{23}o_{134}$ $o_{124}o_{134}$ $o_{125}$ $o_{145}$ $o_{245}$
$o_{123}o_{14}$ $o_{123}o_{234}$ $o_{14}o_{234}$ $o_{15}$ $o_{245}$ $o_{1245}$
$o_{123}o_{24}$ $o_{123}o_{134}$ $o_{24}o_{134}$ $o_{25}$ $o_{145}$ $o_{1245}$
$o_{123}o_{124}$ $o_{123}o_{34}$ $o_{124}o_{34}$ $o_{125}$ $o_{145}$ $o_{245}$
Five of the rows can be derived by rotating the $bc$ indices positively and one negatively, wrapping $1\rightarrow 4$. The other two rows, $o_{2}o_{13}\rightarrow o_{13}o_{4}$ and $o_{12}o_{13}\rightarrow o_{123}o_{4}$, do not follow a pattern nor do the $d$ columns when compared with Table 10, so such pattern recognition is ignored. Since Table 11 contains AAA non-associative cycles this gives another $21\times 4\times 3=252$ zero divisor pairs.
Now all other $bc$ pairs that don’t include $o_{5}$ can be covered. This involves triads with overlapping AAA and BBA cycles but the AAA cases still have four modes and the BBA have prime and extended modes only. In Table 12 the $d$ column provides the AAA structure and the $d_{1}$ and $d_{2}$ columns the BBA structure. The 21 sets of 3-cycles in Table 12 thus represent $21\times 3\times 4=252$ zero divisor pairs in the $d$ column and the $d_{1}$ and $d_{2}$ columns provide $21\times 3\times 2=126$ solutions each which means the table provides $252+126\times 2=504$ zero divisor pairs.
Table 12 can be replicated, apart from the $d$ column, with $o_{4}$ replaced with $o_{5}$ in the $bc$ cycles, as shown in Table 13. The $d_{1}$ and $d_{2}$ columns have the same BBA structure and modes as Table 12 which represents $21\times 3\times 4=252$ solutions. Tables 12 and 13 provide the only B non-associative triads in $d_{1}$ and $d_{2}$ columns giving $21\times 3\times 2\times 2=252$ triads, as expected.
Adding all solutions from Tables 10 through 13 gives $252+252+504+252=1260$ which is enumerated in Table 14 under the ${\mathbb{U}_{2}}$ zero divisors.
Table 13: BBA only Primary Zero Divisors in ${\mathbb{U}_{2}}$
bc d1 d2 bc d1 d2
$o_{1}o_{25}$ $o_{345}$ $o_{2345}$ $o_{3}o_{125}$ $o_{145}$ $o_{245}$
$o_{1}o_{35}$ $o_{245}$ $o_{2345}$ $o_{13}o_{15}$ $o_{245}$ $o_{1245}$
$o_{1}o_{235}$ $o_{245}$ $o_{345}$ $o_{13}o_{25}$ $o_{145}$ $o_{245}$
$o_{2}o_{15}$ $o_{345}$ $o_{1345}$ $o_{13}o_{125}$ $o_{145}$ $o_{245}$
$o_{2}o_{35}$ $o_{145}$ $o_{1345}$ $o_{23}o_{15}$ $o_{245}$ $o_{1245}$
$o_{2}o_{135}$ $o_{145}$ $o_{345}$ $o_{23}o_{25}$ $o_{145}$ $o_{1245}$
$o_{12}o_{15}$ $o_{345}$ $o_{1345}$ $o_{23}o_{125}$ $o_{145}$ $o_{245}$
$o_{12}o_{35}$ $o_{145}$ $o_{1345}$ $o_{123}o_{15}$ $o_{245}$ $o_{1245}$
$o_{12}o_{135}$ $o_{145}$ $o_{345}$ $o_{123}o_{25}$ $o_{145}$ $o_{1245}$
$o_{3}o_{15}$ $o_{245}$ $o_{2345}$ $o_{123}o_{125}$ $o_{145}$ $o_{245}$
$o_{3}o_{25}$ $o_{145}$ $o_{1245}$
Table 14 provides the number of non-associative elements and zero divisors for the ultronions up to level 10. These are labelled ${\mathbb{U}_{m}}$ where $n=m+3$ is the number of generators. Each generator level defines an algebra with $N=2^{n}-1$ pure basis elements and $\binom{N}{3}$ ordered product triples. Removing the associative triads and dividing by 28 provides the number of octonion and pseudo-octonion algebra copies which corresponds with the multiples of 84 zero divisors in the next level ultracomplex algebra. The zero divisors for ${\mathbb{U}_{1}}$ correspond to the AAA silo and those in ${\mathbb{U}_{2}}$ to all of the triads in the AAA and BBA silos. At higher levels, the silos after the first two have overlapping zero divisors. ${\mathbb{U}_{3}}$ has new zero divisors from the ACC and XBB silos that need to be filtered but does not require the BXC silo. ${\mathbb{U}_{4}}$, ${\mathbb{U}_{5}}$, ${\mathbb{U}_{6}}$ and ${\mathbb{U}_{7}}$ add new zero divisors from the BXC silo and non-cycle triads but not the CAB silo. At these higher levels the overlapping cycles that can show 2 or even just 1 mode.
Table 14: Ultronion Algebras Cardinality
Label Pure Basis Non-Associative Zero Divisors
Size Triads 28 Factor Count 84 Factor
${\mathbb{O}}$ 7 28 1 0 0
${\widetilde{\mathbb{O}}}$ 7 28 24
${\mathbb{U}_{1}}\equiv{\mathbb{S}}$ 15 420 15 84 1
${\rm\widetilde{\mathbb{U}}_{1}}$ 15 420 180
${\mathbb{U}_{2}}$ 31 4,340 155 1,260 15
${\mathbb{U}_{3}}$ 63 39,060 1,395 13,020 155
${\mathbb{U}_{4}}$ 127 330,708 11,811 117,180 1,395
${\mathbb{U}_{5}}$ 255 2,720,340 97,155 992,124 11,811
${\mathbb{U}_{6}}$ 513 22,064,980 788,035 8,161,020 97,155
${\mathbb{U}_{7}}$ 1023 177,736,020 6,347,715 66,194,940 788,035
5 Split Cayley-Dickson Algebras
Equation (4) can also be defined with a positive number of $\epsilon_{i}=1$ for $i\in{\mathbb{N}}_{1}^{n}$. Such unitary generators are represented as $u_{i}$ with $u_{i}^{2}=1$ for $i\in{\mathbb{N}}_{1}^{n}$. This provides a mixed sign basis with elements $o_{\alpha}u_{\beta}$ with $\alpha,\beta\in{\mathrm{B}(n)}$ such that $\alpha\cap\beta=\emptyset$. Lemma 3 shows that squaring such elements is $-1$ for even length of $\beta$ and is $+1$ for an odd length. This is now applied to each level of the graded algebras.
By the binomial theorem, the number of basis elements is the sum over all grades including the number one, as shown in (1). An enumeration of the binomial theorem is expressed in Pascal’s triangle that starts as
1
1 1
1 2 1
1 3 3 1
1 4 6 4 1
and is described by the recurrence relation that each position after the first is the sum of the two elements above. For graded algebras the triangle enumerates Cayley-Dickson algebras starting with the reals, ${\mathbb{R}}$. Each row after this is the statement that the number of terms of any $k$-grade is given by the number of terms at grade $k$ without the generator $o_{n}$ or $u_{n}$ plus those that include this generator from grade $k-1$. Pascal’s triangle is horizontally symmetric because for each term of grade $k$, the complimentary term with grade $n-k$ contains the same $k$ indices removed from the $n$ generator case. For $n$ odd there are an equal number of terms with $o_{n}$ or $u_{n}$ as without in the complimentary terms. For $n$ even the same applies apart from the middle grade $k=n/2$. But $\binom{n}{n/2}$ is always even for $n$ even and by the definition of combinations it distributes the $n$ generators equally into an even set, half of which contain $o_{n}$ or $u_{n}$ and the others exclude it. Of course, the left-hand side has grade $k=0$ that specifies the basis number 1 while the right-hand side, the compliment of 1, has $k=n$ and always contains $o_{n}$ or $u_{n}$.
Definition.
The pure trace of the multiplication table for the algebra with $n$ generators is the sum of the squares of the $12^{n}-$ 1 pure basis terms, which is $-(2^{n}-1)$ for Cayley-Dickson algebra at level $n$. Thus for octonions this is $-7$.
Theorem 12 (Trace theorem).
The pure trace for all of the split ultracomplex algebras is one, as it is for split octonions, ${\widetilde{\mathbb{O}}}$, split quaternions, ${\widetilde{\mathbb{H}}}$, and Lorentz numbers, ${\widetilde{\mathbb{C}}}$.
Proof.
For any blade $o_{\alpha}u_{\beta}$, $\alpha$ and $\beta$ from disjoint sets of ${\mathrm{B}(n)}$ then Lemma 3 shows that $(o_{\alpha}u_{\beta})^{2}=\pm 1$, $-1$ if $\beta$ has even length and $+1$ if $\beta$ has odd length. Starting at $n=1$ the Lorentz numbers, $\widetilde{\mathbb{C}}$, have one blade, $u_{1}$, so has pure trace one. Split-quaternions, $\widetilde{\mathbb{H}}$, can have pure basis, $\{o_{1},u_{2},o_{1}u_{2}\}$ or $\{u_{1},u_{2},u_{12}\}$, both of which satisfy the theorem. By induction, the basis at level $n-1$ is assumed to have pure trace one and either $o_{n}$ or $u_{n}$ will multiplying all basis elements. By the binomial theorem recurrence relation the existing pure basis is doubled by multiplying all existing basis elements by $o_{n}$ and adding the $o_{n}$ (multiplication by 1). Extending each $\alpha$ in this way does not change the parity of $\beta$ so this doubles the pure trace count and adds the -1 trace part for $o_{n}$, keeping the pure trace to one.
Now applying the same recurrence relation with $u_{n}$ swaps even and odd parities of each $\beta$ in the new multiplied terms when excluding the unit, 1. Hence a trace of one at level $n-1$ has terms added with opposite trace parity giving a total parity of zero. But the extra term, $u_{n}$, from multiplication by 1, regains the pure trace of one. ∎
Theorem 13 (Split Cayley-Dickson algebras).
All Cayley-Dickson algebras with any positive number of $\epsilon_{i}=-1$ for $i\in{\mathbb{N}}_{1}^{n}$ are isomorphic to each other so may be represented as the algebra of unitary generators, $u_{1},u_{2},\cdots,u_{n}$, called the split algebra, either $\widetilde{\mathbb{C}}$, $\widetilde{\mathbb{H}}$, $\widetilde{\mathbb{O}}$ or ${\rm\widetilde{\mathbb{U}}_{m}}$, $m=n-3>0$.
Proof.
From Theorem 1, all blades $x,y$ anti-commute, $xy=-yx$, if $x\neq y$, and by Lemma 3 have squares $x^{2}=\pm 1$, positive for an odd number of unitary generators and negative otherwise. A basis of the algebra is $x=o_{\alpha}u_{\beta}$, such that $\alpha\cap\beta=\emptyset$, $\forall\alpha,\beta\in{\mathrm{B}(n)}$. By Theorem 12, any positive number of unitary generators produce a basis with pure trace one. Selecting a set of $n$ of $\alpha+\beta$ so that $(o_{\alpha}u_{\beta})^{2}=1$ but the set elements are not all from the split algebra at level $n-1$ then from Lemma 1 this set generates the whole of ${\mathrm{B}(n)}$ and the corresponding basis elements can be relabelled as new generators $u_{1}^{\prime},u_{2}^{\prime},\ldots,u_{n}^{\prime}$. ∎
The unitary Type 1, Type 2 and Type 3 associativity scheme for triads does not change from the non-split case because the same sign changes apply to each associativity term. This means the non-associativity types A, B, C and X are also appropriate but the zero divisors do change. These start appearing for the split octonions, $\widetilde{\mathbb{O}}$, which have $12$ zero divisor pairs which, of course, have X-non-associativity. They have dual partners so only six need to be shown with a partial cyclic structure,
$(u_{1},u_{2},u_{13}),(u_{1},u_{12},u_{3}),(u_{2},u_{12},u_{3}),$
$(u_{1},u_{12},u_{123}),(u_{2},u_{12},u_{123}),(u_{3},u_{13},u_{123})\text{.}$
Split-octonions have idempotents of the form $(1+u_{\alpha})$ with nilpotents $(1+u_{\alpha})(1-u_{\alpha})$ for $\alpha\in{\mathbb{N}}_{1}^{3}\cup\{123\}$. These are derived from the four positive square blades of $\widetilde{\mathbb{O}}$, $u_{\alpha}^{2}=1$ with $u_{123}^{2}=1$. Each idempotent can be multiplied on the left by the three blades that do not contain $u_{\alpha}$, which form a left ideal and gives $16$ nilponents. Multiplying by different terms of the nilpotent gives another zero divisor. For example, $(1+u_{1})$ has $(-u_{2}+u_{12})(u_{3}+u_{13})=0$, which satisfies the definition for a zero divisor since $u_{12}u_{3}u_{13}=-u_{2}$. Hence ideals overlap with zero divisors for split algebras.
Table 15: Split Sedenion Zero Divisors
b c d blades b c d blades
$u_{1}$ $u_{2}$ $u_{13},u_{14},-u_{134}$ $u_{12}$ $u_{3}$ $u_{34},-u_{134}$
$u_{1}$ $u_{3}$ $u_{23},u_{14},u_{124}$ $u_{12}$ $u_{13}$ $-u_{4},u_{14}$
$u_{1}$ $u_{23}$ $u_{4},u_{24}$ $u_{12}$ $u_{4}$ $u_{34}$
$u_{1}$ $u_{4}$ $u_{24},u_{34},u_{1234}$ $u_{12}$ $u_{14}$ $-u_{134}$
$u_{1}$ $u_{24}$ $-u_{134},-u_{234}$ $u_{3}$ $u_{4}$ $u_{14},u_{24}$
$u_{1}$ $u_{34}$ $-u_{234}$ $u_{3}$ $u_{14}$ $-u_{124}$
$u_{2}$ $u_{3}$ $u_{13},u_{24},-u_{124}$ $u_{3}$ $u_{24}$ $-u_{124}$
$u_{2}$ $u_{13}$ $u_{4},u_{14}$ $u_{13}$ $u_{4}$ $u_{24}$
$u_{2}$ $u_{4}$ $u_{14},u_{34},u_{1234}$ $u_{13}$ $u_{14}$ $-u_{124}$
$u_{2}$ $u_{14}$ $-u_{134},-u_{234}$ $u_{23}$ $u_{4}$ $u_{14}$
$u_{2}$ $u_{34}$ $u_{134}$ $u_{23}$ $u_{24}$ $-u_{124}$
Split-sedenions, ${\rm\widetilde{\mathbb{U}}_{1}}$, have 180 zero divisor pairs derived from 108 X, 48 A and 24 C-non-associative triads which includes extending the zero divisor to allow for unitary products. Unlike the split octonions, ${\rm\widetilde{\mathbb{U}}_{1}}$ has a full cyclic structure with correspondence of the $d$ terms for all cycle triads. The X types have only two modes and not necessarily the dual. The A and C types either have only the primary mode or all four modes with A type having three such cases and C have just two. This reduces to the 39 primaries shown in Table 15 and with 21 modes then the 3 triad cycles for each triad gives the 180 zero divisor pairs for ${\rm\widetilde{\mathbb{U}}_{1}}$. The $d$ blades shown as negative in the table require $(-a+b)(c+d)$ for all first cycles but some second and third cycles need the extended dual mode as well. The two exceptions are $(u_{1},u_{2},-u_{134})$ and $(u_{2},u_{3},-u_{124})$ which need the extended dual mode with the standard positive zero divisor.
An overlap between split and non-split zero divisors does not occur. The triad $(u_{12},u_{13},u_{14})$ and its product, $u_{12}u_{13}u_{14}=-u_{1234}$, contains terms that are all imaginary, which should be analogous to the non-split triad $(o_{12},o_{13},o_{14})$ but this has product $+o_{1234}$ providing a zero divisor with opposite parity.
6 Summary
The Cayley-Dickson algebras have been assumed to be symmetric in the permutation group but this is only valid for hypercomplex numbers. The octonions introduce non-associativity but in a symmetric form. Using a graded notation provides a rigorous definition to the order of basis elements and exposes the asymmetry of the power-associative Cayley-Dickson algebras. The octonion non-associative triads are alternate, which means swapping basis elements in the associator changes its sign. The term ultracomplex numbers has been introduced to cover the case of the associator toggling between zero and non-zero when swapping elements of the associator. This extends the symmetric non-associativity to another three types that exhibit asymmetric properties in cycles that leads to multiples of 84 zero divisor pairs and modes that, along with cycles, can reduce these down to 7, in some cases. The total number of zero divisors for each algebra is also derived.
The graded approach also applies to split ultracomplex numbers by showing that, like the hypercomplex case there is only one split algebra at each level of the Cayley-Dickson construction. In this case it is convenient to use $\epsilon=-1$ for all generators which then have positive square $u_{i}^{2}=1$. The convenient symmetry thereafter is that all odd length grades are unitary and all even grades are imaginary. The non-associativity from the non-split case carries over for identical indices with identical zero divisors for the imaginary terms whereas the unitary zero divisors can be related to ideals of the algebra or imaginary zero divisors. Open questions for future work are to establish connections of split zero divisors to ideals, gain further understand how zero divisors affect Malcev algebras, investigate the relationships between the pseudo-octonion algebras and establish why only three of the six are contained in power-associative algebras.
The author gratefully acknowledges the invaluable assistance and encouragement of James Chapell, Jacqui Ramagge and Derek Abbott for improving terminology, mathematical understanding and ongoing assistance with the publishing of this article. The author also acknowledges the support of an Australian Government Research Training Program Scholarship. All of this work was verified with the use of an octonion/ultronion calculator written in Python. The github URL for the calculator is